### v1

# GARLIC: GAussian Representation LearnIng for spaCe partitioning

Abstract
We introduce GARLIC (GAussian Representation LearnIng for spaCe partitioning), a novel indexing structure based on $N$-dimensional Gaussians for efficiently learning high-dimensional vector spaces. Our approach is inspired from Gaussian splatting techniques, typically used in 3D rendering, which we adapt for high-dimensional search and classification. We optimize Gaussian parameters using information-theoretic objectives that balance coverage, assignment confidence, and structural and semantic consistency. A key contribution is to progressively refine the representation through split and clone operations, handling hundreds of dimensions, thus handling varying data densities. GARLIC offers the fast building times of traditional space partitioning methods (e.g., under $\sim 5$ min build time for SIFT1M) while achieving $\sim 50\%$ Recall10@10 in low-candidate regimes. Experimental results on standard benchmarks demonstrate our method’s consistency in (a) $k$-NN retrieval, outperforming methods, such as Faiss-IVF, in fast-recall by using about half their probes for the same Recall10@10 in Fashion-MNIST, and (b) in classification tasks, beating by $\sim 15\%$ accuracy other majority voting methods. Further, we show strong generalization capabilities, maintaining high accuracy even with downsampled training data: using just $1\%$ of the training data returns $\sim 45\%$ Recall@1, thus making GARLIC quite powerful for applications requiring both speed and accuracy.
1 Introduction
Let $X=\{x_{i}\}_{i=1}^{n}\subset\mathbb{R}^{d}$ be a set of high-dimensional embeddings, derived from handcrafted descriptors (e.g., SIFT1M [37]) or learned via neural networks, which encode semantic and geometric similarity. Such embeddings are commonly used in tasks such as $k$-nearest neighbor ($k$-NN) retrieval and classification. In retrieval, the goal is to return the $k$ closest points in $X$ to a query $q\in\mathbb{R}^{d}$, i.e., $\mathcal{N}_{k}(q)=\arg\min_{\begin{subarray}{c}S\subset X,|S|=k\end{subarray}}\sum_{x\in S}\|q-x\|^{2}$, typically under Euclidean or cosine distance [2]. For classification, each point $x\in X$ carries a label $y_{x}\in\mathcal{Y}$, and the query’s label is predicted via majority vote among neighbors: $\hat{y}(q)=\arg\max_{c\in\mathcal{Y}}\sum_{x\in\mathcal{N}_{k}(q)}\mathbb{I}[y_{x}=c]$, where $\mathbb{I}[\cdot]$ is the indicator function [12]. Alternatively, a learned decision function $f:\mathbb{R}^{d}\to\mathcal{Y}$ can be trained directly over the embedding space [23]. In both cases, the density, structure, and discriminative power of $X$ are crucial to performance.
Approaches that build a static index, such as hashing techniques [25], and vector quantization methods [19], which partition a set of data points $X$, construct the index independently of the data distribution. Rather than fitting the indexing structure to the spatial and geometric properties of set of data point $X$, these methods project the data points to a pre-defined and fixed structure. Such methods usually impose generic assumptions on the distribution of $X$, to provide theoretical error bound guarantees. Partitions based on such assumptions often misalign with underlying high-dimensional data distributions, which is also known as the curse of clustering, while distance metrics fall short due to curse of dimensionality.
Several ANN methods build indices aligned with the data distribution, including LSH variants [14, 5], quantization-based approaches [19, 15], clustering methods [6], and tree-based partitioning [31]. These techniques aim to accelerate retrieval by organizing the space into regions aligned with approximate neighborhood structure. While effective in certain regimes, they exhibit key limitations: high training cost for large-scale data, fragile performance under distribution shift, and rigid, non-adaptive structure post-training. In contrast, deep learning methods model the data manifold directly, capturing local density and semantic continuity to enable flexible, unified retrieval and classification [9, 45].
This begs the question: How can we efficiently learn the high dimensional space to perform fast, accurate and semantic approximate nearest neighbor searches? In other words can we learn a geometric representation $\mathcal{G}_{\theta}$ of $\mathbb{R}^{d}$, parameterized by, e.g., Gaussians or neural fields, that enables efficient, robust solutions to both retrieval and classification, while adapting to the structure and distribution of $X$?
1.1 Related Work
Traditional ANNS approaches. Approximate nearest neighbor search (ANNS) methods traditionally fall into three categories: (i) Indexing, including LSH-based hashing [25, 19, 46], randomized projections and Johnson–Lindenstrauss-based embeddings [4], vector quantization [20, 26, 3], and tree-based methods [28]; (ii) Sketching, which compresses data via hash functions to accelerate approximate distance computations [16, 35, 36, 34, 13, 27, 43]; and (iii) Graph-based search, which performs greedy traversal on structures such as Navigating Spreading-out Graphs [18] or Hierarchical Navigable Small World Graphs [38]. While effective in practice, these methods typically rely on fixed partitions or hand-crafted heuristics that do not adapt to the geometry of learned embedding spaces.
The rise of high-dimensional embeddings from deep and geometric representation learning has shifted the focus of ANNS toward learning-based approaches. In contrast to classical features, these embeddings are structured: they reflect semantic similarity, lie on low-dimensional manifolds, and exhibit highly nonuniform density. This calls for indexing methods that do not impose rigid, predefined partitions, but instead learn representations that align with the intrinsic geometry of the data.
Learning-based indexing. A prominent direction is learning to index, where models are trained to build adaptive partitions over the data [14, 21]. Neural LSH [14], for example, learns a partitioning of $\mathbb{R}^{d}$ via a classifier trained on a balanced $k$-NN graph. However, such balance constraints often conflict with real data distributions—especially under long-tailed or clustered regimes—leading to poor locality. Beyond hard partitioning, recent approaches have explored classification-based routing [24] and neural scoring of clusters [11], aiming to improve query-time ranking. Yet these methods still rely on discretized, often rigid space decompositions. In contrast, we advocate for a geometric learning perspective: learning smooth, structure-aware representations that are density and semantic aligned. Such representations enable unified treatment of retrieval and classification, generalize under distribution shift.
Contributions. We propose GARLIC, a method for learning geometry-aware representations of high-dimensional spaces via a set of Gaussians optimized for both approximate nearest neighbor retrieval and classification. GARLIC learns a probabilistic representation that adapts to local data density and semantic similarity, aligning the shape and placement of Gaussians with the underlying structure of the data. In particular, our method adapts to regions with varying intrinsic dimensionality (e.g., curves, surfaces, volumes), enabling each Gaussian to act as an informative geometric primitive and local manifold approximator, rather than a static partitioning cell.
  • We introduce a geometric representation learning approach inspired by Gaussian Splatting [30], that implicitly partitions the input space, optimized for both retrieval and classification using a unified training objective.
  • We develop adaptive refinement strategies (split, clone) tailored for high-dimensional spaces, and a localized quantization method for handling density variation, improving search efficiency and robustness w.r.t. noise.
  • We design a novel information-theoretic objective combining Mahalanobis-based divergence, entropy-reducing confidence, and geometry-preserving anchoring, aligning the Gaussian structure with the data manifold.
  • We benchmark GARLIC for retrieval and classification. Our approach exhibits state-of-the-art performance on both tasks w.r.t. runtime-recall tradeoff, especially for small candidate sets. The source code of our approach will be publicly available.
2 Method
GARLIC uses a set of $d$-dimensional Gaussians $\mathcal{G}=\{\mathcal{N}(\mu_{i},\Sigma_{i})\}_{i}$, that adapt to the underlying data distribution, provided as a set of points $x\in X\subset\mathbb{R}^{d}$, to represent high-dimensional space for efficient ANNS and classification. We choose Gaussians because they capture both local geometry and statistical structure, unlike simpler volumes such as boxes or zonotopes, and express anisotropy, curvature, and density variation in a unified, continuous form. The closed form Mahalanobis scoring and differentiability also make them well-suited for end-to-end optimization. Starting with data points, we represent the space with the set of Gaussians $\mathcal{G}$ parameterized by means and covariance matrices via Cholesky decomposition (Sec. 2.1). We optimize the coefficients (mean and cholesky) of the Gaussians using objectives that aim to balance coverage, point assignment confidence, and structural consistency (Sec. 2.2). Unlike traditional partitioning methods that enforce or tend towards balanced divisions [14, 19, 6, 31, 1, 41, 39], our approach adaptively refines the set of Gaussians $\mathcal{G}$ through splitting and cloning operations, creating a data-dependent partitioning that accommodates varying densities (Sec. 2.3). To further enhance search efficiency, i.e., narrow down the number of searches, a local hyper-spherical quantization index within each Gaussian bucket (Sec. 2.4) is built. At query time, our method uses Mahalanobis distances for fast bucket selection and prioritized bin search for efficient retrieval (Sec. 2.5). This combination of techniques allows GARLIC to effectively capture the natural clustering structure of high-dimensional data while maintaining search performance, as illustrated in Fig. 1.
Figure 1: Overview of GARLIC. The method starts with input vector points that are used to initialize a set of Gaussians parameterized by mean $\boldsymbol{\mu}$ and covariance matrices via Cholesky decomposition. During optimization, these Gaussians are refined through our novel objectives that balance coverage, assignment confidence, and structural consistency. The representation is adaptively improved through split and clone operations creating a data-dependent semantic partitioning that better reflects varying data densities in different regions of the feature space.
2.1 Gaussian Parameterization
Each Gaussian $g_{i}$ is defined by its mean $\boldsymbol{\mu_{i}}\in\mathbb{R}^{d}$ and covariance matrix $\boldsymbol{\Sigma}_{i}\in\mathbb{R}^{d\times d}$, parameterized as $\boldsymbol{\Sigma}_{i}=\mathbf{L}_{i}{\mathbf{L}_{i}}^{\top}$, $\mathbf{L}_{i}\in\mathbb{R}^{d\times d}$ being a lower triangular Cholesky factor of matrix $\boldsymbol{\Sigma}_{i}$, ensuring positive semi-definiteness (a requirement for valid Gaussian distributions). The means $\boldsymbol{\mu}_{i}$ are initialized using Κ-Μeans++ [7] on a subset of the training set. Cholesky factors $\mathbf{L}_{i}$ are initialized as:
$\mathbf{L}_{i}=\log(\bar{\delta_{\mathcal{E}}})\cdot\mathbf{I}_{d}+\boldsymbol{\epsilon},$
where $\mathbf{I}_{d}\in\mathbb{R}^{d\times d}$ is the identity matrix, and $\bar{\delta_{\mathcal{E}}}$ is the mean Euclidean distance from each $\boldsymbol{\mu}_{i}$ to its three nearest neighbors. This ensures that the initial scale of each Gaussian reflects the local data density. The perturbation $\boldsymbol{\epsilon}\in\mathbb{R}^{d\times d}$ is a random lower-triangular matrix, defined below:
$\boldsymbol{\epsilon}_{jk}=\begin{cases}2\cdot\sigma(r_{jk})-1,&j>k\\ 0,&\text{otherwise}\end{cases}$
where $r_{jk}\sim\mathcal{U}(0,0.01)$ and $\sigma(\cdot)$ is the sigmoid function. This results in diagonal-dominant $\mathbf{L}_{i}$, which stabilizes the optimization process while enabling flexible, anisotropic covariance structures. This initialization strategy attempts to align the initial Gaussian scales with the local data density and to ensure sufficient diversity among Gaussians for adaptive refinement in later stages.
2.2 Optimization Objective
We adapt ideas from 3D Gaussian Splatting (3DGS) [29] to high-dimensional nearest neighbor search. While 3DGS performs well in 3D scene modeling, its reliance on Structure-from-Motion (SfM) initialization and view-space heuristics breaks down in high dimensions, where geometry is abstract and dimensions encode heterogeneous features. Splitting and cloning based on spatial cues become unreliable, and optimization suffers in sparse regions. These issues are exacerbated by the curse of dimensionality and clustering, which degrade distance metrics and make balanced partitioning ineffective. High-dimensional indexing demands adaptive, data-aware refinement, something 3DGS heuristics cannot provide.
To handle these issues, inspired by Self-Supervised learning techniques [9, 45], that use information theoretic objectives to increase the capacity of the representation space, we introduce similar information objectives to guide optimization towards a more informative and balanced solution without excessive prior usage. More specifically, we introduce a divergence-based objective that acts as a reconstruction loss, and regularize it to prevent information explosion (i.e., uncontrolled growth and excessive overlap of Gaussians). The divergence loss $\mathcal{L}_{\text{div}}$ is defined as:
$\mathcal{L}_{\text{div}}=\frac{1}{N}\sum_{\mathbf{x}\in X}\left(\min_{g_{i}\in\mathcal{G}}\delta_{M}(\mathbf{x},g_{i})-\tau\right)^{+},$
where $\delta_{M}(\mathbf{x},g_{i})=||\mathbf{L}_{i}^{-1}(\mathbf{x}-\boldsymbol{\mu}_{i})||_{2}$ denotes the Mahalanobis distance, $(\cdot)^{+}=\max(0,\cdot)$ and $\tau$ is a standard deviation threshold. It penalizes points that fall outside a Gaussian’s coverage radius ($\delta_{M}(\mathbf{x}_{i},g_{i})>\tau$), encouraging Gaussians to expand and cover these points, while points inside ($\delta_{M}(\mathbf{x}_{i},g_{i})\leq\tau$) do not contribute, allowing controlled expansion but preventing infinite growth.
Still, Gaussians can overlap, leading to redundant information. To surpass this issue, and mitigate fuzzy assignments of points to Gaussians, we introduce a covariance-based regularization, which encourages each Gaussian to dominate its assigned points. Specifically, given a point $\mathbf{x}$, we define its coverage set, i.e., the set of Gaussians that satisfy the coverage radius constraint, as $\mathcal{M}(\mathbf{x})=\{g_{i}\in\mathcal{G}\mid\delta_{M}(\mathbf{x},g_{i})\leq\tau\}.$
Then, we compute the normalized soft-assignment probabilities based on Euclidean distances ($\delta_{\mathcal{E}}$) as $p_{i}(\mathbf{x})=e^{-\delta_{\mathcal{E}}(\mathbf{x},\boldsymbol{\mu}_{i})}/\sum_{g_{j}\in\mathcal{M}(\mathbf{x})}e^{-\delta_{\mathcal{E}}(\mathbf{x},\boldsymbol{\mu}_{j})}+\epsilon,\ \forall g_{i}\in\mathcal{M}(\mathbf{x})$. The covariance loss $\mathcal{L}_{\text{cov}}$ is defined as:
$\mathcal{L}_{\text{cov}}=1-\frac{1}{N}\sum_{\mathbf{x}\in X}\max_{g_{i}\in\mathcal{M}(\mathbf{x})}p_{i}(\mathbf{x})$
This loss encourages highly confident (low-entropy) assignments, thereby reducing ambiguity and stabilizing optimization. To prevent excessive expansion of Gaussians caused by the divergence objective and ensure that each Gaussian accurately aligns with its assigned points, we introduce the anchor loss $\mathcal{L}_{\text{anchor}}$:
$\mathcal{L}_{\text{anchor}}=\frac{1}{d|G|}\sum_{g_{i}\in G}\left(||\boldsymbol{\mu}_{i}-\hat{\boldsymbol{\mu}}_{i}||_{2}^{2}+\alpha||\mathbf{L}_{i}{\mathbf{L}_{i}}^{\top}-\hat{\boldsymbol{\Sigma}}_{i}||_{F}^{2}\right),$
where $\hat{\boldsymbol{\mu}}_{i}$ and $\hat{\boldsymbol{\Sigma}}_{i}=\mathrm{Cov}(x\in B_{g_{i}})$ are the empirical mean and covariance of points assigned to Gaussian $g_{i}$, and $\alpha$ is a hyperparameter balancing position and shape. This loss constrains Gaussians by anchoring them closely to their local data distributions, counteracting uncontrolled growth from other loss terms and maintaining geometric alignment and interpretability. Finally, our loss function is defined as:
$\mathcal{L}=\lambda_{\text{div}}\cdot\mathcal{L}_{\text{div}}+\lambda_{\text{cov}}\cdot\mathcal{L}_{\text{cov}}+\lambda_{\text{anchor}}\cdot\mathcal{L}_{\text{anchor}},$
where $\lambda_{\text{div}},\lambda_{\text{cov}}\text{ and }\lambda_{\text{anchor}}$ are hyperparameters balancing the importance of each term.
(a) Clone operation. (light red) region where $d_{M}(x,g_{i})\leq\tau$; (blue) outer shell $\tau\leq d_{M}(x,g_{i})\leq e\tau$; (gray) $\{x:\ d_{M}(x,g_{i})\geq e\tau\}$; (red) mean of the new gaussian; (green) new covariance matrix. (purple - lime) denote increasing data density.
(b) Grad flow: ${\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\mathcal{L}_{\text{div}}}$, ${\color[rgb]{0.72,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.72,0,0}\pgfsys@color@cmyk@stroke{0}{0.89}{0.94}{0.28}\pgfsys@color@cmyk@fill{0}{0.89}{0.94}{0.28}\mathcal{L}_{\text{cov}}}$, ${\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}\mathcal{L}_{\text{anchor}}}$
Figure 2: Gaussian refinement and associated loss gradients.
2.3 Adaptive Refinement
Controlling the number of Gaussians during training is key for adapting to data complexity; allocating more in dense regions and fewer in simpler ones. We employ an adaptive refinement strategy that adjusts the Gaussian set based on local point density. Prior approaches rely on positional gradients to guide refinement [29], but these become sparse and unreliable in high dimensions. In our case, we also require search efficiency, so we refine Gaussians whose bucket cardinality exceeds a threshold: $|\mathcal{B}_{g}|>\gamma\cdot|X|$. Splitting is performed by applying DBSCAN [17] to detect two dense clusters within the bucket. If unsuccessful, we fall back to K-Means to produce new means $\boldsymbol{\mu}_{1},\boldsymbol{\mu}_{2}$, while covariances are scaled down as $\mathbf{L}_{1}=\mathbf{L}_{2}=\alpha\cdot\mathbf{L}$, with $\alpha<1$.
For cloning, we target Gaussians with a high ratio of outside points to interior ones $\mathcal{B}_{g_{i}}^{in}=\{\mathbf{x}\in X:\delta_{M}(\mathbf{x},g_{i})\leq\tau\}$. Instead of using all nearest outside points, we focus on a boundary region, and select $k=\rho\cdot|\mathcal{B}_{g}^{out}|$ points, where $\rho\in(0,1)$ is a sampling ratio and $\mathcal{B}_{g_{i}}^{out}=\{\mathbf{x}\in X:\tau<\delta_{M}(\mathbf{x},g_{i})<e\tau\text{ and }g_{i}=\arg\min_{g_{j}\in G}\delta_{M}(\mathbf{x},g_{j})\}$ with $e>1$ is the set of all boundary points. The sampled subset $S=\{\mathbf{x}_{1},\mathbf{x}_{2},...,\mathbf{x}_{k}\}\subset\mathcal{B}_{g}^{out}$ is chosen randomly without replacement. This sampling approach serves several purposes, reducing computational cost, as density estimation in high dimensions is expensive, providing statistical robustness by focusing on representative points and helping to avoid outliers that might exist in the boundary region as Fig. 2 showcases.
The candidate Gaussians are picked with regards to the ration of boundary to interior points: $|\mathcal{B}_{g}^{out}|/|\mathcal{B}_{g}^{in}|>\beta$. The center of the new Gaussian is placed at the point with the highest local density in the boundary region, such that for a point $\mathbf{p}$ in the boundary region, we compute its local density as:
$\rho(\mathbf{p})=\frac{1}{\frac{1}{k}\sum_{j=1}^{k}\delta_{\mathcal{E}}(\mathbf{p},\text{NN}_{j}(\mathbf{p}))},$
here $\text{NN}_{j}(\mathbf{p})$ is the $j$-th nearest neighbor of $\mathbf{p}$ and $\delta_{\mathcal{E}}(\cdot,\cdot)$ is the Euclidean distance. We select the point with the highest density: $\mathbf{p}^{*}=\arg\max_{\mathbf{p}}\rho(\mathbf{p})$. The covariance matrix is then cloned, with $\mathbf{L}_{new}=\mathbf{L}$.
2.4 Quantization
After optimizing our Gaussians, we assign points that fall outside the coverage radius to the nearest Gaussian such that our final buckets are:
$B_{g_{i}}=\{\mathbf{x}\in X\mid\delta_{M}(\mathbf{x},g_{i})\leq\tau\}\cup\{\mathbf{x}\in X\mid\delta_{M}(\mathbf{x},g_{i})>\tau\text{ and }g_{i}=\arg\min_{g_{j}\in G}\delta_{M}(\mathbf{x},g_{j})\}$
Algorithm 1 Build Index
1:Input: Dataset $\mathcal{D}$ of $d$-dimensional points
2:Initialize Gaussians on $\mathcal{D}$
3:$\mathcal{G}\leftarrow$ Obtain $K$ Gaussians $g_{i}=\mathcal{N}(\mu_{i},L_{i})$ (Sec. 2.22.3)
4:for each point $p$ do
5:    Assign $p$ to all Gaussians where $d_{M}(p,g_{i})<\tau$, $\forall i\in[K]$
6:    if $p$ not assigned to any Gaussian then
7:         Assign $p$ to nearest Gaussian w.r.t. $d_{M}$
8:    end if
9:end for
10:for each Gaussian $g_{i}\in\mathcal{G}$ do
11:    $\mathcal{Q}_{i}\leftarrow$ Quantize indexed data of $g_{i}$ (Sec. 2.4)
12:end for
13:return $\{(g_{i},\mathcal{Q}_{i})\}_{i\in[K]}$
While our Gaussian-based buckets provide an effective organization of the data space, exhaustive search within buckets becomes prohibitively expensive as dimensionality increases. For high-dimensional nearest neighbor search, we must avoid calculating distances to all points within candidate buckets. Standard approximate search methods, however, often do not account for the unique structure provided by our Gaussian-based partitioning.
We introduce a quantization scheme that leverages the inherent properties of our Gaussian model. The key insight is that once we have assigned points to appropriate Gaussian buckets, we can apply efficient Euclidean-based quantization within each bucket’s local coordinate system, reducing the number of distance computations required at query time.
Algorithm 2 Query Index
1:Input: Index structure set $\{(g_{i},\mathcal{Q}_{i})\}_{i\in[K]}$, query $q$
2:$g_{i}\leftarrow\arg\min_{i:g_{i}\in\mathcal{G}}d_{M}(q,g_{i})$
3:$B\leftarrow$ Traverse subset of bins from $\mathcal{Q}_{i}$ (Sec. 2.4)
4:Measure $L_{2}$ distance to candidates in $B$
5:Terminate if condition (Sec. 2.5)
6:return nearest neighbor of $q$ in $B$
For each bucket $B_{g}$, we apply PCA to reduce dimensionality while preserving the local structure $\mathbf{P}_{g}=\text{PCA}(\{\mathbf{x}-\bar{\mathbf{x}}_{g}\mid\mathbf{x}\in B_{g}\},r)$ where $\bar{\mathbf{x}}_{g}$ is the mean of points in bucket $B_{g}$, $r$ is the reduced dimensionality (typically $r\ll d$, constant in practice), and $\mathbf{P}_{g}\in\mathbb{R}^{d\times r}$ contains the principal components. Each point $\mathbf{x}\in B_{g}$ is projected into this lower-dimensional space $\mathbf{x}^{r}=\mathbf{P}_{g}^{T}(\mathbf{x}-\bar{\mathbf{x}}_{g})$, after which its transformed to hyperspherical coordinates $\mathbf{s}=\text{cart2sph}(\mathbf{x}^{r})$, where the transformation for an $r$-dimensional point yields coordinates $\mathbf{s}=(s_{1},s_{2},\ldots,s_{r})$.
The hyperspherical space is then partitioned into bins defined by $n_{r}$ radial divisions and $n_{a}$ angular divisions per dimension: $\mathcal{B}_{g}=\{B_{i,\mathbf{j}}\mid i\in\{1,\ldots,n_{r}\},\mathbf{j}\in\{1,\ldots,n_{a}\}^{r-1}\}$ with boundaries $r_{i}=r_{\min}+(r_{\max}-r_{\min})\cdot i/n_{r},\ i\in\{0,\ldots,n_{r}\}$ and $\theta_{j,k}=\theta_{\min,k}+(\theta_{\max,k}-\theta_{\min,k})\cdot j/n_{a},\ j\in\{0,\ldots,n_{a}\},k\in\{1,\ldots,r-1\}$. Alg. 1 illustrates the construction process of the index structure.
2.5 Inference
At query time, for a query point $\mathbf{q}$, we employ a multi-stage approach to efficiently retrieve approximate nearest neighbors. First, we identify the $k$ nearest Gaussians using the Mahalanobis distance. For each selected Gaussian bucket $g$, we project the query into the bucket’s local PCA space, $\mathbf{q}^{r}_{g}=\mathbf{P}_{g}^{T}(\mathbf{q}-\bar{\mathbf{x}}_{g})$, and convert it to hyperspherical coordinates $\mathbf{s}_{g}=\text{cart2sph}(\mathbf{q}^{r}_{g})$. We then compute the distances to all bins in the bucket and sort them in ascending order. For each bin $B_{i,\mathbf{j}}$ with boundaries defined by $r$-intervals $[r_{i},r_{i+1}]$ and angular intervals $[\theta_{j,k},\theta_{j+1,k}]$, we compute the distance $d(\mathbf{s}_{g},B_{i,\mathbf{j}})=\min_{\mathbf{y}\in B_{i,\mathbf{j}}}\|\mathbf{s}_{g}-\mathbf{y}\|_{2}.$
This minimization problem is efficiently solved by first checking if $\mathbf{s}_{g}$ is inside the bin. If it is outside, we solve the constrained optimization problem by initializing a point $\mathbf{y}^{0}$ with its radial component clipped to $[r_{\min},r_{\max}]$ and angular components set to the midpoints of their respective intervals. We then minimize $\|\mathbf{s}_{g}-\mathbf{y}\|_{2}$ subject to the bin boundary constraints using the L-BFGS-B [10] algorithm, which respects the bounds $r_{\min}\leq y_{1}\leq r_{\max}$ and $\theta_{\min,j}\leq y_{j+1}\leq\theta_{\max,j}$ for the radial and angular components respectively.
We then search for points in bins sequentially according to their sorted distances, computing exact Euclidean distances between the query and the points in each bin. The search terminates under the condition that we have examined a predefined percentage of bins (typically 30%). Given a query, Alg. 2 showcases the procedure of retrieving the nearest neighbor using our index structure.
3 Experiments
Datasets. We evaluate our method on three benchmark datasets: SIFT1M [37] (128-dimensional image descriptors with one million points), MNIST [32] and Fashion-MNIST [44] (784-dimensional vectors from 28×28 grayscale images). These datasets provide diverse characteristics in terms of extrinsic and intrinsic dimensionality, scale, and data distribution, offering a comprehensive evaluation framework in Euclidean space. Additional information about the datasets and their licenses are provided in the Supplementary Material (SM).
OursK-MeansN-LSHITQCP-LSH
HP-LSHPCA TreeFaiss-IVFFaiss-IVFPQFS
Ours 10%Ours 5%Ours 2.5%Ours 1%
(a) Efficiency
(b) MNIST
(c) Fashion-MNIST
(d) SIFT1M
(e) SIFT1M
(f) MNIST
(g) Fashion-MNIST
[…]
Evaluation. We evaluate our method on retrieval and classification tasks across the aforementioned datasets. For retrieval, we report Recall10@10 and Recall@1 versus the number of retrieved candidates (probes), reflecting the trade-off between accuracy and efficiency. Recall10@10 measures how many of the true top-10 neighbors are in the retrieved set, while Recall@1 checks if the top-ranked result is the true nearest neighbor. The candidate count controls query time and serves as the main efficiency metric. For MNIST and Fashion-MNIST, we assess classification via majority voting over retrieved candidates: $\hat{y}(q)=\arg\max_{c}\sum_{x\in C_{q}}\mathbb{I}[\text{label}(x)=c]$. This evaluates how well the partitioning preserves semantic similarity.
Baselines and Comparisons. We evaluate our proposed method against several state-of-the-art approximate nearest neighbor search algorithms such as K-Means [6], Neural-LSH (N-LSH) [14], Hyperplane LSH (HP-LSH) [25, 5], Cross-polytope LSH (CP-LSH) [5], ITQ [19], PCA Tree [31, 1, 41, 39], and Faiss-IVF and Faiss-IVFPQFS [15]. For classification tasks, we additionally compare against standard $k$-Nearest Neighbors ($k$-NN) [12, 42], which performs exhaustive search for exact neighbors, and $k$-K-Nearest Neighbors ($kK$-NN) [33], which adjusts the neighborhood size based on local curvature estimates. Each method was evaluated under identical conditions, with hyperparameters optimized according to their original publications to ensure fair and meaningful comparisons across both nearest neighbor retrieval and classification performance.
Table 1: Classification accuracy via majority voting. Ours-1: $\arg\min_{g}\delta_{M}(q,g)$; Ours-2: $\delta_{M}(q,g)\leq\tau$; Ours-3: top-$k$ Gaussians. We highlight the best, second-best and third-best results.
Dataset
Method MNIST $\uparrow$ Fashion-MNIST $\uparrow$
k-NN 0.7576 0.7133
kK-NN 0.7609 0.7796
K-Means 0.9118 0.7589
PCA Tree 0.7750 0.7427
ITQ 0.7156 0.6459
HP-LSH 0.5483 0.6643
CP-LSH 0.7801 0.7128
N-LSH 0.4142 0.4307
[…]
Ours-$3$ 0.9175 0.7879
Results and Analysis. We present performance curves across datasets, reporting Recall10@10 and Recall@1 as a function of the number of retrieved candidates. As shown in Fig. 3, our method consistently outperforms state-of-the-art baselines across both metrics. Notably, it exhibits a significant advantage in the low-probe regime: for example, on SIFT1M, our method achieves Recall@1 of $\sim 69\%$ using $\sim 3400$ candidates, compared to 61% for Faiss-IVF and 55% for Faiss-IVFPQFS using similar candidate counts. This efficiency gap is most evident below the 80% recall threshold, where traditional methods struggle to match the accuracy without a sharp rise in probe cost. Fig. 3(a) highlights the efficiency trade-off between training time and retrieval quality, measured via Recall@1-per-probe. All our variants lie in the top-left region of the plot, offering options ranging from ultra-fast lightweight models (e.g., Ours 1%) to more accurate, higher-capacity variants (e.g., Ours 10%). In contrast, competing methods tend to occupy lower-efficiency or higher-cost zones, demonstrating a less favorable balance between indexing overhead and retrieval quality.
(a) Generalization
(b) Time-recall@1
(c) Quantization
(d) $|\mathcal{G}|$ impact
Figure 4: (a) Generalization with limited training data. (b) Time-recall@1 tradeoff. (c) Quantization: efficient search via partial bucket scanning. (d) Impact of initial number of Gaussians.
Furthermore, for classification tasks on MNIST and Fashion-MNIST, our approach demonstrates superior classification accuracy compared to the other benchmarks, shown in Tab.1. This enhancement is consistent with theoretical expectations, as our methodology learns the inherent probabilistic structure of the training data rather than depending on predetermined heuristics, static partitioning, or hash-based approximations. By dynamically adapting to the statistical properties of the data distribution, our method identifies more semantically coherent neighbors. Our results suggest that GARLIC acts as a local manifold modeler, not just an indexing method offering smoother, distribution-aware retrieval and classification.
Table 2: Effect of loss terms and adaptive refinement on average Recall10@10 / Probe (×$\times 10^{5}$). Higher is better.
Configuration Performance $\uparrow$
Loss terms
   w/ all 10.43
   $\mathcal{L}_{div}+\mathcal{L}_{cov}$ 9.82
   $\mathcal{L}_{div}+\mathcal{L}_{anchor}$ 7.87
   $\mathcal{L}_{div}$ 7.17
[…]
   w/ both 10.43
   w/ split 8.65
   w/ clone 8.47
   w/o any 2.79
Ablation study. We investigate the robustness and data efficiency of our method through controlled downsampling of the training set. As shown in Fig. 4(a), our method consistently maintains strong performance across varying training sizes. Notably, even with only $1\%$ of the training data, our model outperforms the Faiss-IVF variants of up to $10\%$ training size in terms of Recall@1 versus the number of retrieved candidates. This trend remains consistent across higher percentages, demonstrating that our approach learns a compact yet highly effective representation of the data distribution. Furthermore, Fig. 4(b) highlights the favorable scaling of our method with respect to training time, while Fig. 4(c) demonstrates the impact of our quantization scheme. In particular, it exhibits strong performance across a broad quantization range ($20-100\%$), with only the most aggressive setting (10%) leading to degradation. This indicates that our quantization strategy is robust and well-aligned with our model structure. Fig. 4(c) shows that increasing the number of initial Gaussians accelerates convergence. However, the primary role of adaptive refinement is not to fix the final count, but to control their spatial density.
Tab. 2 analyzes the impact of loss terms and adaptive refinement. Split and clone are pivotal for performance, while the covariance term ($\mathcal{L}_{\text{cov}}$) significantly boosts retrieval. The anchor term ($\mathcal{L}_{\text{anchor}}$) has minor effect on raw performance but is essential for making Gaussians geometrically informative. Additional ablations and results, covering hyperparameter sensitivity, design variants, and the ability of Gaussians to capture non-linear, informative structure, are provided in the SM.
Limitations. GARLIC uses Mahalanobis distance, which assumes Euclidean structure and is not directly compatible with angular similarity. This can be addressed by replacing Gaussians with angular analogues such as von Mises–Fisher distributions or by embedding into hyperspherical spaces. Although GARLIC performs strongly in low-probe regimes, achieving very high recall (e.g., $\geq 95\%$) requires more computation. This recall-latency trade-off can be mitigated by increasing the number of Gaussians and structuring them with trees (e.g., KD-tree or Ball-tree over Gaussian means), reducing query cost. The space complexity scales quadratically with the dimension ($O(Kd^{2})$) due to full anisotropic covariances; however, this can be alleviated by using diagonal or low-rank approximations, or by quantizing the Cholesky factors. Lastly, while the method is mildly sensitive to initialization, its impact is reduced by the information-theoretic objectives and the progressive refinement through split and clone operations.
4 Conclusions
We introduced GARLIC, a geometric structure that learns the underlying distribution for both approximate nearest neighbor search and classification. By combining information-theoretic objectives with adaptive refinement (split and clone), and representing the space via anisotropic Gaussians with local quantization, our method achieves strong performance in both retrieval and classification, particularly in low-probe regimes. Experiments demonstrate competitive recall-efficiency tradeoffs and robustness under severe data reduction, highlighting its generalization capabilities.
Acknowledgments
PR and IE were partially supported by project MIS 5154714 of the National Recovery and Resilience Plan Greece 2.0 funded by the European Union under the NextGenerationEU Program. PR, CT, IC, GI and IE have been partially supported by the European Union and the “Greece 2.0” national recovery and resilience plan, under the call "RESEARCH-CREATE-INNOVATE" (Code: ΤΑΕΔΚ-06168).
Appendix A Supplementary Material
This supplementary material provides additional details and experiments as mentioned in the main paper. Sec. A.1 discusses and provides detailed computational time and space complexity for the build and query procedures of our proposed method. Sec. A.2 contains additional technical details regarding the datasets used and the experimental setup, as well as the training pipeline. Sec. A.3 provides ablation studies to validate our choices regarding the tuned hyperparameters of our method and the components of it (discussed in Sec.4 of the main paper). Sec. A.4 contains visualization of diagnostics and statistics regarding our proposed method, in order to provide a comprehensive understanding of its behavior.
A.1 Complexity Analysis
We analyze the computational time and space complexity of our method in three parts: index construction, query execution, and storage. The analysis is expressed in terms of standard parameters, including the dataset size $|X|$, embedding dimension $d$, the number of Gaussians $K$, and the reduced PCA dimension $r\ll d$. Our goal is to ensure that each component remains scalable with respect to high-dimensional data and large-scale datasets. We summarize the complexity of each phase below.
Index build complexity.
Let $|X|$ be the number of data points, $d$ the embedding dimension, $K^{\prime}$ the initial number of Gaussians, and $r\ll d$ the reduced dimension used for local quantization. Let $I$ be the number of optimization steps. We denote by $S$, $C$, and $P$ the number of splits, clones, and pruned Gaussians, respectively, and define the final number of Gaussians as $K=K^{\prime}+S+C-P$. Let $|B_{g}|$ be the average bucket size, $c$ the number of K-Means clusters used during splitting, and $k^{\prime}$ the number of candidate points sampled per clone.
For the initialization, since we use K-Means++ on $K^{\prime}$ total cluster centers, we need $\mathcal{O}(K^{\prime}\cdot d\cdot|X|)$ time. For the optimization part, we need to perform a total number of $I$ iterations of full Mahalanobis-based point-to-Gaussian assignment, thus a total of $\mathcal{O}(I\cdot|X|\cdot K\cdot d^{2})$ worst-case time. Separate from the optimization, we analyze the split, clone and prune operations that are not applied on every iteration of the optimization. (i) The split operation runs DBSCAN or K-Means (with $c$ clusters), thus for a total of $S$ such operations we would need $\mathcal{O}(S\cdot(|B_{g}|\cdot d\cdot c+d^{2}))$; (ii) the clone operation locates the subset of points outside the Gaussian’s boundary (e.g., between $\tau\cdot\sigma$ and $(1+\epsilon)\cdot\tau\cdot\sigma$) for which it identifies new local modes, thus for a total of $C$ operations, it leads to $\mathcal{O}(C\cdot(k^{\prime 2}\cdot d+d^{2}))$; and (iii) the prune operation simply removes low-cardinality Gaussians and reassigns the points to the nearest active Gaussian, which takes $\mathcal{O}(P\cdot|B_{g}|\cdot d)$ time. For the quantization of each Gaussian, PCA is performed on all points inside the Gaussian, which projects the data into reduced local bases. In total, for the quantization we need $\mathcal{O}(K^{\prime}\cdot|B_{g}|\cdot d^{2}+|X|\cdot d\cdot r)$ time.
Query complexity.
Let $K$ be the number of Gaussians, $k$ the number selected per query, $d$ the dimension, $r$ the PCA dimension, $b$ the number of bins per Gaussian, $T$ the number of optimization steps to find the shortest distance from the query point to the boundary of a spherical bin in the reduced PCA space, $\rho$ the probed bin ratio, and $\beta$ the average bin size. For a single query, we first need to measure distances from the set of Gaussians, which takes $\mathcal{O}(K\cdot d^{2})$. Then, for the $k$ nearest Gaussians we need to locate the subset of data to be examined. For this, for each of the selected $k$ Gaussians, we need to compute the local PCA projections ($\mathcal{O}(k\cdot d\cdot r)$), then compute and sort the spherical distances to all $b$ bins ($\mathcal{O}(k\cdot b\cdot r\cdot T)$), of which only the $\rho$ fraction is probed. From each, up to $\beta$ candidates are gathered and re-ranked using Euclidean distance, which needs $\mathcal{O}(k\cdot\rho\cdot b\cdot\beta\cdot d)$ time. In practice, the re-ranking factor dominates the complexity, which is sublinear.
Space complexity.
Let $K$ be the number of Gaussians, $d$ the data dimension, and $N=|X|$ the dataset size. The model stores mean vectors $\boldsymbol{\mu}\in\mathbb{R}^{K\times d}$, Cholesky parameters $\mathbf{L}\in\mathbb{R}^{K\times d\times d}$, and Buckets storing point indices, requiring $\mathcal{O}(N)$ space. Thus, the total space complexity is:
$\mathcal{O}\left(K\cdot(d^{2}+d)+N\right)$
where $K\cdot d^{2}$ dominates. Still, space complexity can be reduced to $\mathcal{O}\left(K\cdot d\right)$ by enforcing diagonal covariance matrices, at the expense of reduced expressiveness in anisotropic regions of the space.
A.2 Implementation details
This subsection summarizes all implementation and training-specific parameters used in our model, including optimizer schedules, architectural constants, and adaptive procedures such as splitting, cloning and pruning. These details provide context for reproducibility and support the complexity analysis in the main paper. Furthermore, there are dataset specifics, such as licenses and descriptions.
Table 3: Dataset Information
Dataset Dimension Size Description License
SIFT1M 128 1M Image descriptors CC0
MNIST 784 70K Handwritten digits CC BY-SA 3.0
Fashion-MNIST 784 70K Fashion items MIT
Datasets.
We evaluate our method on three standard benchmark datasets: (1) SIFT1M [37], containing one million 128-dimensional SIFT descriptors that capture scale- and rotation- invariant local image features; (2) MNIST [32], consisting of 70,000 grayscale handwritten digit images (28×28 pixels) flattened to 784-dimensional vectors; and (3) Fashion-MNIST [44], a more challenging variant with the same format but featuring 10 categories of fashion items. For retrieval tasks, we used the ANN-Benchmark [8] versions of these datasets (available at http://ann-benchmarks.com/sift-128-euclidean.hdf5, http://ann-benchmarks.com/mnist-784-euclidean.hdf5, and http://ann-benchmarks.com/fashion-mnist-784-euclidean.hdf5) to ensure a standardized comparison with existing methods. For MNIST and Fashion-MNIST classification experiments, we utilized the PyTorch torchvision implementations to maintain consistency with established classification benchmarks.
Experimental setup.
Our method was implemented in Python, using optimized libraries such as PyTorch [40] and NumPy [22]. The experiments were carried out on an Intel Core i7-7820X CPU (16 threads), a Quadro RTX 8000 GPU (48 GB VRAM) and 125 GiB of RAM.
Training configuration.
Batch sizes of $20000$ are used for SIFT, and $5000$ for MNIST and Fashion-MNIST datasets, with normalization by subtracting the mean and dividing by the standard deviation. The number of training epochs is $250$, with a typical early stop at $120$. Gaussian updates are scheduled with two phases: a warm-up phase lasting $35$ epochs and an optimization phase where structural operations like splitting and cloning are triggered every $35$ epochs and pruning every $60$.
Learning rate schedule.
The learning rates follow a linear warm-up and exponential decay scheme. Specifically, for the Cholesky parameters, the rates are subjected to a warm-up phase from $1\times 10^{-7}$ to $5\times 10^{-4}$, followed by a decay to $9\times 10^{-5}$. In terms of the means, they are warmed up from $1\times 10^{-7}$ to $9\times 10^{-3}$, and subsequently decay to $3\times 10^{-3}$. Notably, the learning rate associated with the means is maintained at a relatively higher level than that of the covariances. This approach is designed to promote the adjustment of the Gaussian centers rather than the expansion of their radii.
Adaptive refinement.
Splitting is applied to Gaussians with cardinality exceeding a fraction of the dataset $\gamma=1\times 10^{-2}$, using DBSCAN or K-Means, as fallback, with $c=2$ clusters. The covariance of each new Gaussian is scaled down by $9\times 10^{-1}$. Cloning selects dense regions just outside the Gaussian boundary, defined by a Mahalanobis shell with inner threshold $\tau$ and outer threshold $(1+\epsilon)\cdot\tau$, where $\epsilon=2.2$ controls the shell thickness. From the set of points referenced in $(\tau,e\tau]$, a random selection of $60\%$ is made. Cloning is not performed on a Gaussian unless its cardinality surpasses a threshold specified by $8\times 10^{-4}|X|$. Gaussians that have degenerated into a single point are eliminated. Pruning is executed at intervals of every $60$ epochs.
Quantization.
Local PCA is performed per bucket using top-$3$ eigenvectors. Reduced points are quantized using a spherical grid with $n_{\text{radial}}=6$ and $n_{\text{angular}}=4$, forming directional bins per Gaussian.
Loss.
The total loss is a weighted sum of three components: divergence $\lambda_{div}=1.0$, covariance $\lambda_{cov}=1.0$, and anchor term $\lambda_{anchor}=10^{-2}$, with a weight $\alpha=10^{-1}$ that balances position and shape. When calculating the $\mathcal{L}_{cov}$ loss, a numerical epsilon of $1\times 10^{-12}$ is used to ensure stability.
A.3 Additional Ablation Studies
[…]
(a) Loss terms
(b) Split & clone impact
(c) ReLU parameters $\tau$
(d) High-recall
[…]
Architectural components.
In Fig. 5(5(a)), the contribution of each loss term to the optimization process is analyzed. The results indicate that the covariance loss $\mathcal{L}_{\text{cov}}$ produces the most significant performance improvement, while the anchor loss $\mathcal{L}_{\text{anchor}}$ functions as an effective regularizer, grounding each Gaussian by aligning it with its corresponding point distribution. Fig. 5(5(b)) demonstrates the effects of our adaptive refinement operations, where the split and clone mechanisms improve retrieval performance when used together, each contributing by providing complementary benefits to the quality of representation. Fig. 5(5(c)), addresses the parameter $\tau$ within $\mathcal{L}_{div}$, demonstrating that $\tau=3$ is the optimal choice, particularly in scenarios involving a limited number of probes. Values surpassing those presented in our study ($\tau>3$) result in less optimal outcomes. For instance, $\tau=4$ demonstrates a Recall10@10 value of 0.670.67~{}0.670.67 when evaluated with $65000$ probes.
A way to address the limitation associated with high-recall scenarios, as discussed in the limitations paragraph of the main paper, is show in Fig. 5(5(d)). Specifically, it involves assessing all Gaussian candidates that meet the condition $d_{M}(x,g)\leq\tau$, rather than solely the one that minimizes the Mahalanobis distance. This approach is also semantically justified, considering that multiple Gaussians may occupy the same region. Consequently, a hybrid approach can be used, utilizing the $\arg\min$ strategy for scenarios with a low number of probes and employing all candidates that satisfy condition $d_{M}(x,g)\leq\tau$ for high-recall cases, thus effectively addressing the limitation.
(a) Initialization
(b) Split threshold $\gamma$
(c) Clone threshold $e$
(d) Dimensionality effect
[…]
Table 4: Effect of anisotropy on average Recall10@10 / Probe (×$\times 10^{5}$). Higher is better. As anisotropy decreases, performance degrades due to excessive probe usage.
Configuration Performance $\uparrow$
Covariance structure
   Anisotropic 10.43
   Diagonal 11.04
   Isotropic 10.63
Validation of hyperparameters.
Fig. 6 provides a systematic evaluation of our key hyperparameters. In Fig. 6(6(a)), we investigate initialization strategies for Gaussian centers, revealing that K-Means initialization consistently outperforms alternatives. Although farthest point sampling (FPS) achieves comparable recall, it requires approximately 50% more probes, and catching up to K-Means requires more training time. In Fig. 6(6(b)), the split threshold $\gamma$, controlling the density condition for Gaussian subdivision, is examined. The optimal values are $\gamma=0.005$ and $\gamma=0.01$, though $\gamma=0.005$ requires more time due to more frequent splits. As for the clone threshold $e$, depicted in Fig. 6(6(c)), which determines the outer shell for cloning, it remains stable across different values, with $e=2.2$ being preferable in situations requiring high recall. Finally, Fig. 6(6(d)) examines dimensionality effects using Fashion-MNIST data resized to various dimensions. The results confirm that GARLIC maintains strong performance across a wide range of dimensionalities, highlighting its robustness to the feature space complexity.
Covariance structure vs performance.
We conduct a targeted study to assess the impact of different covariance configurations-namely, full (anisotropic), diagonal, and isotropic-baffect the performance of GARLIC. As shown in Table 4, reducing the Gaussian expressiveness from full to diagonal and then to isotropic leads to a notable decline in average Recall10@10 per probe. While diagonal and isotropic configurations reduce both the parameter count and computation per Gaussian to a linear level, $\mathcal{O}(k\cdot d)$, they lead to aggressive pruning and an increased number of probes to make up for the representational loss. This suggests that space complexity cannot be drastically reduced without harming retrieval quality because simpler Gaussian parameterizations result in degraded locality and coverage.
A.4 Qualitative Results
To gain a more comprehensive understanding of the model’s behavior, we present a collection of diagnostic visualizations applied to the Fashion-MNIST, MNIST, and SIFT datasets. These plots illustrate structural characteristics, including local coverage, reconstruction fidelity, density patterns, and curvature statistics of the learned Gaussian components. All visualizations are conducted on a randomly sampled subset of training points, utilizing the learned parameters independently of test data supervision.
Figure 7: Diagnostic visualizations across Fashion-MNIST (left), MNIST (middle), and SIFT (right). Top row: minimum Mahalanobis reconstruction error; middle row: average Gaussian coverage per point; bottom row: relationship between local density and Gaussian coverage. Bottom row’s colormap depicts frequency-density, in logarithmic scale.
Coverage and reconstruction diagnostics.
Figure 7 illustrates three diagnostic views, each calculated for a different dataset (Fashion-MNIST, MNIST, and SIFT), to evaluate the accuracy of the Gaussian models in representing the datasets. The top row includes the reconstruction landscape, which visualizes the minimum Mahalanobis distance from each point to any Gaussian. For each dataset, we employ PCA to project the points onto a two-dimensional space and calculate the average reconstruction error of proximate points on a grid. This heatmap highlights how well the Gaussian shells approximate the distribution of data throughout the space.
The middle row shows the coverage landscape, which counts how many Gaussians fall within the Mahalanobis threshold $\tau$ for each data point. Coverage is computed per point, projected to 2D, and smoothed via k-NN averaging over a grid. This plot reflects the redundancy and spatial spread of the Gaussian coverage. We observe that areas with low reconstruction error tend to be have high coverage.
The bottom row depicts the relationship between local density and Gaussian coverage. For each point, we compute its local density via the inverse mean distance to its 10 nearest neighbors and correlate this with its coverage count. The resulting 2D histograms reveal structural patterns where regions of higher density generally exhibit greater coverage while sparse regions receive fewer assignments. This aligns with our goal of achieving balanced coverage while maintaining good reconstruction fidelity.
Figure 8: Curvature diagnostics across Fashion-MNIST (left), MNIST (middle), and SIFT (right). Row 1: standard deviation of local curvature vs radius; Row 2: mean local curvature vs radius; Row 3: Gaussian coverage vs curvature std; Row 4: Gaussian coverage vs curvature mean. Coloring accounts for density.
Curvature-based diagnostics.
Fig. 8 provides four views exploring the relationship between curvature and structural properties of the Gaussian assignments across datasets, to examine wheather the learned model is informative and geometrically consistent with the data $x\in X$.
The first row shows the standard deviation of local curvature as a function of the average radius ($l_{2}$) per Gaussian. For each Gaussian, we collect nearby assigned points (under Mahalanobis threshold $\tau$), compute their curvature via PCA-based local flatness, and report the standard deviation. Each bin aggregates Gaussians by radius and variation in curvature, highlighting the stability of their local geometry, where for each dataset curvatures deviations tend to be around $\sim 0.06$.
The second row reports the mean curvature of each Gaussian against its average radius. This indicates the intrinsic dimensionality and shape complexity of regions assigned to Gaussians of different spatial extent. In general, we see that mean curvature tends to be $\sim 0.5$, suggesting a moderate level of local non-linearity, especially for Gaussians with smaller support. As the radius increases, curvature remains relatively stable, indicating consistent local geometry across scales.
The third row depicts how Gaussian coverage (number of assigned points) varies with the curvature standard deviation of the assigned region. We observe that most Gaussians exhibit low curvature variability (std $\sim 0.06$), indicating that points within each Gaussian tend to have similar geometric structure. Moreover, there is no clear correlation between coverage and curvature std, implying that heavily used Gaussians are not more geometrically diverse than others. This suggests a form of balanced representation capacity across Gaussians.
The fourth row shows coverage as a function of mean curvature, where Gaussians with high coverage have curvature patterns similar to those with low coverage.
Figure 9: Spectral diagnostics across datasets (SIFT, MNIST, Fashion-MNIST; columns) and Gaussian properties (rows). Each subplot shows a different measure: anisotropy, band energy, diagonality, effective dimension, and explained variance.
Fig. 9 demonstrates the capability of GARLIC to capture the intrinsic geometric structure of high-dimensional data with locally varying dimensionality. The anisotropy histograms (first row) reveal how Gaussians adapt to regions of different local dimensionality, with $\lambda_{\text{max}}/\lambda_{\text{min}}$ ratios ranging from nearly isotropic to highly stretched configurations across all datasets. Rather than simply partitioning space uniformly, Gaussians adapt their shapes to the underlying manifold structure, as confirmed by the diagonality histograms (third row) and effective dimension measurements (fourth row). This representation allows us to model stratified data where different intrinsic dimensionalities coexist, allowing the data to become pancake-like for surface regions, needle-like for curve regions, and ball-like for volumetric areas. Unlike traditional space partitioning methods, GARLIC models the underlying data distribution probabilistically, not just approximating distances for retrieval. The explained variance surfaces (bottom row) show that Gaussians efficiently capture the relevant dimensions at each location, enabling estimation of true manifold distances rather than just Euclidean distances to samples. This provides more semantically meaningful results in regions where the intrinsic dimensionality is lower than the ambient space. This completes the supplementary material.

### v2

@@
 Abstract
-We introduce GARLIC (GAussian Representation LearnIng for spaCe partitioning), a novel indexing structure based on $N$-dimensional Gaussians for efficiently learning high-dimensional vector spaces.
-Our approach is inspired from Gaussian splatting techniques, typically used in 3D rendering, which we adapt for high-dimensional search and classification.
-We optimize Gaussian parameters using information-theoretic objectives that balance coverage, assignment confidence, and structural and semantic consistency.
-A key contribution is to progressively refine the representation through split and clone operations, handling hundreds of dimensions, thus handling varying data densities.
-GARLIC offers the fast building times of traditional space partitioning methods (e.g., under $\sim 5$ min build time for SIFT1M) while achieving $\sim 50\%$ Recall10@10 in low-candidate regimes.
-Experimental results on standard benchmarks demonstrate our method’s consistency in (a) $k$-NN retrieval, outperforming methods, such as Faiss-IVF, in fast-recall by using about half their probes for the same Recall10@10 in Fashion-MNIST, and (b) in classification tasks, beating by $\sim 15\%$ accuracy other majority voting methods.
-Further, we show strong generalization capabilities, maintaining high accuracy even with downsampled training data: using just $1\%$ of the training data returns $\sim 45\%$ Recall@1, thus making GARLIC quite powerful for applications requiring both speed and accuracy.
+We present GARLIC, a representation learning approach for Euclidean approximate nearest neighbor (ANN) search in high dimensions.
+Existing partitions tend to rely on isotropic cells, fixed global resolution, or balanced constraints, which fragment dense regions and merge unrelated points in sparse ones, thereby increasing the candidate count when probing only a few cells.
+Our method instead partitions $\mathbb{R}^{d}$ into anisotropic Gaussian cells whose shapes align with local geometry and sizes adapt to data density.
+Information-theoretic objectives balance coverage, overlap, and geometric alignment, while split/clone refinement introduces Gaussians only where needed.
+At query time, Mahalanobis distance selects relevant cells and localized quantization prunes candidates.
+This yields partitions that reduce cross-cell neighbor splits and candidate counts under small probe budgets, while remaining robust even when trained on only a small fraction of the dataset.
+Overall, GARLIC introduces a geometry-aware space-partitioning paradigm that combines information-theoretic objectives with adaptive density refinement, offering competitive recall–efficiency trade-offs for Euclidean ANN search.
 1 Introduction
-Let $X=\{x_{i}\}_{i=1}^{n}\subset\mathbb{R}^{d}$ be a set of high-dimensional embeddings, derived from handcrafted descriptors (e.g., SIFT1M [37]) or learned via neural networks, which encode semantic and geometric similarity.
-Such embeddings are commonly used in tasks such as $k$-nearest neighbor ($k$-NN) retrieval and classification.
-In retrieval, the goal is to return the $k$ closest points in $X$ to a query $q\in\mathbb{R}^{d}$, i.e., $\mathcal{N}_{k}(q)=\arg\min_{\begin{subarray}{c}S\subset X,|S|=k\end{subarray}}\sum_{x\in S}\|q-x\|^{2}$, typically under Euclidean or cosine distance [2].
-For classification, each point $x\in X$ carries a label $y_{x}\in\mathcal{Y}$, and the query’s label is predicted via majority vote among neighbors: $\hat{y}(q)=\arg\max_{c\in\mathcal{Y}}\sum_{x\in\mathcal{N}_{k}(q)}\mathbb{I}[y_{x}=c]$, where $\mathbb{I}[\cdot]$ is the indicator function [12].
-Alternatively, a learned decision function $f:\mathbb{R}^{d}\to\mathcal{Y}$ can be trained directly over the embedding space [23].
-In both cases, the density, structure, and discriminative power of $X$ are crucial to performance.
-Approaches that build a static index, such as hashing techniques [25], and vector quantization methods [19], which partition a set of data points $X$, construct the index independently of the data distribution.
-Rather than fitting the indexing structure to the spatial and geometric properties of set of data point $X$, these methods project the data points to a pre-defined and fixed structure.
-Such methods usually impose generic assumptions on the distribution of $X$, to provide theoretical error bound guarantees.
-Partitions based on such assumptions often misalign with underlying high-dimensional data distributions, which is also known as the curse of clustering, while distance metrics fall short due to curse of dimensionality.
-Several ANN methods build indices aligned with the data distribution, including LSH variants [14, 5], quantization-based approaches [19, 15], clustering methods [6], and tree-based partitioning [31].
-These techniques aim to accelerate retrieval by organizing the space into regions aligned with approximate neighborhood structure.
-While effective in certain regimes, they exhibit key limitations: high training cost for large-scale data, fragile performance under distribution shift, and rigid, non-adaptive structure post-training.
-In contrast, deep learning methods model the data manifold directly, capturing local density and semantic continuity to enable flexible, unified retrieval and classification [9, 45].
-This begs the question: How can we efficiently learn the high dimensional space to perform fast, accurate and semantic approximate nearest neighbor searches?
-In other words can we learn a geometric representation $\mathcal{G}_{\theta}$ of $\mathbb{R}^{d}$, parameterized by, e.g., Gaussians or neural fields, that enables efficient, robust solutions to both retrieval and classification, while adapting to the structure and distribution of $X$?
+Let $X=\{x_{i}\}_{i=1}^{n}\subset\mathbb{R}^{d}$ be a finite point set, $q\in\mathbb{R}^{d}$ a query, and $\delta_{\mathcal{E}}:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}_{\geq 0}$ the Euclidean distance.
+For an integer $k\geq 1$, the exact $k$-Nearest Neighbor Search (NNS) problem returns the $k$ closest points $N_{k}(q)\subseteq X$.
+Its approximate variant, $k$-ANN, relaxes this by requiring $A_{k}(q)\subseteq X$, $|A_{k}(q)|=k$, such that $\max_{a\in A_{k}(q)}\delta_{\mathcal{E}}(q,a)\leq c\,\delta_{\mathcal{E}}\!\big(q,x_{(k)}(q)\big),$ for some approximation factor $c\geq 1$, where $x_{(k)}(q)$ denotes the $k$-th true neighbor of $q$.
+We restrict attention to Euclidean spaces of hundreds of dimensions, and to indices defined by partitions of $\mathbb{R}^{d}$ into cells, where a query inspects only a few cells and re-ranks the resulting candidates.
+Nearest neighbor search in this setting is a canonical problem of high-dimensional geometry and algorithms, with consequences across information retrieval, computer vision, robotics, and data analysis (Lowe, 2004, Cai et al., 2021, Shakhnarovich et al., 2008, Aumüller et al., 2020, Douze et al., 2024).
+ANN algorithms attempt to reduce cost in two independent ways.
+Sketch-based techniques (Razenshteyn, 2017, Wang et al., 2014) attempt to compress every point into a short-coded representation, a summary, so that approximate distances can be quickly evaluated.
+Index-based methods (Gani et al., 2016) pre-partition the space, and examine only a subset of the point set, at query time.
+The two approaches are complementary and often combined in practice.
+The focus of this work is on indexing methods, and most specifically space-partitioning in the $\bigl(\mathbb{R}^{d},\delta_{\mathcal{E}}\bigr)$ metric space (Aumüller et al., 2020).
+The space is divided into cells $\mathcal{B}_{g}$, each storing data points, and a query touches only those stored in the cell that contains the query point (plus a few neighbors for higher recall).
+Space-partition indices are practical and efficient with their small space overhead, each cell stores a representative, e.g., a centroid, and a list of point IDs, far less than e.g., graph indices need (Malkov & Yashunin, 2018).
+Cells can be queried in parallel by different cores, read in one-shot by GPUs, and fetched as one block (I/O call) from the disk storage (Johnson et al., 2019, Douze et al., 2024, Jayaram Subramanya et al., 2019).
+These strengths hold only when the cells are well built, whether by fixed, data-independent rules or by partitions learned from the data.
+What drives performance is how cells are built.
+Data-independent (Andoni & Indyk, 2008, Andoni et al., 2018) rules fix splits a priori (e.g., random hyperplanes, simple trees), so they build in $\mathcal{O}(|X|d)$ time but ignore the geometry of $X$, and recall degrades on clustered or curved regions.
+Data-dependent schemes (e.g., k-means/IVF families (Jegou et al., 2010), learned hashing (Wang et al., 2015)) fit cell parameters to $X$ and usually improve recall per number of candidates visited.
+In practice, partitions often tend to be isotropic, for example Voronoi cells around $k$-means centroids (Lloyd, 1982, Arthur & Vassilvitskii, 2007), and a single global number of cells $K$ is chosen.
+These design choices then cause predictable errors on heavy-tailed data (Clauset et al., 2009).
+Partition resolution, a single global number of cells $K$ applied everywhere, means dense regions get fragmented into many small (near-spherical) cells, while sparse regions are covered by a few large ones (Du et al., 1999).
+Balanced partitions on non-uniform data create a complementary problem: in dense zones they split true neighbors across cells, and in sparse zones they pack unrelated points together to meet the size target (Malinen & Fränti, 2014, Aumüller et al., 2020).
+Neighborhoods are approximately Euclidean only locally; large cells merge unrelated regions, while overly small ones fragment continuous neighborhoods.
+To recover locality one must either increase $K$ or probe (touch), many adjacent cells (Johnson et al., 2019, Lv et al., 2007), which raises candidate (distance) counts and hurts low-probe regimes.
+This leaves a concrete gap: partitions that capture local geometry and adapt their local resolution (effective cell size / expected cell cardinality) to density under a principled objective that balances reconstruction fidelity against candidate count.
+Contributions.
+We propose GARLIC, a geometry-aware space-partition index for Euclidean ANN, optimized under an information-theoretic objective that balances coverage, overlap, and budget efficiency.
+Under this objective, GARLIC learns a probabilistic partition of $\mathbb{R}^{d}$ into Gaussian cells whose shape and placement align with local principal directions and whose sizes adapt to local density, adding capacity only where needed through local adaptive refinement.
+The resulting partition reduces cross-cell neighbor splits and candidate counts under small probe budgets.
+  • Anisotropic, density-adaptive partition.
+GARLIC represents $\mathbb{R}^{d}$ with Gaussian cells that follow local geometry and adapt to density, improving within-cell neighbor cohesion under small candidate budgets.
+(Section 2.1 – 2.3)
+  • Information-theoretic objective.
+We balance coverage, overlap, and probe efficiency via expected Mahalanobis coverage, an assignment-entropy penalty, and geometric anchoring regularization.
+(Section 2.2)
+  • Local adaptive refinement.
+We add Gaussians only where needed through split/clone operations triggered by cell size and spill ratio, avoiding a single global resolution (one $K$ everywhere).
+(Section 2.3)
+  • Budget-centric evaluation and analysis.
+We report competitive performance across multiple accuracy and distortion metrics under candidate and distance budgets on standard Euclidean benchmarks, and provide build/query/space complexity together with ablations that isolate each component’s contribution (Section 3, Appendix A.1).
 1.1 Related Work
-Traditional ANNS approaches.
-Approximate nearest neighbor search (ANNS) methods traditionally fall into three categories: (i) Indexing, including LSH-based hashing [25, 19, 46], randomized projections and Johnson–Lindenstrauss-based embeddings [4], vector quantization [20, 26, 3], and tree-based methods [28]; (ii) Sketching, which compresses data via hash functions to accelerate approximate distance computations [16, 35, 36, 34, 13, 27, 43]; and (iii) Graph-based search, which performs greedy traversal on structures such as Navigating Spreading-out Graphs [18] or Hierarchical Navigable Small World Graphs [38].
-While effective in practice, these methods typically rely on fixed partitions or hand-crafted heuristics that do not adapt to the geometry of learned embedding spaces.
-The rise of high-dimensional embeddings from deep and geometric representation learning has shifted the focus of ANNS toward learning-based approaches.
-In contrast to classical features, these embeddings are structured: they reflect semantic similarity, lie on low-dimensional manifolds, and exhibit highly nonuniform density.
-This calls for indexing methods that do not impose rigid, predefined partitions, but instead learn representations that align with the intrinsic geometry of the data.
-Learning-based indexing.
-A prominent direction is learning to index, where models are trained to build adaptive partitions over the data [14, 21].
-Neural LSH [14], for example, learns a partitioning of $\mathbb{R}^{d}$ via a classifier trained on a balanced $k$-NN graph.
-However, such balance constraints often conflict with real data distributions—especially under long-tailed or clustered regimes—leading to poor locality.
-Beyond hard partitioning, recent approaches have explored classification-based routing [24] and neural scoring of clusters [11], aiming to improve query-time ranking.
-Yet these methods still rely on discretized, often rigid space decompositions.
-In contrast, we advocate for a geometric learning perspective: learning smooth, structure-aware representations that are density and semantic aligned.
-Such representations enable unified treatment of retrieval and classification, generalize under distribution shift.
-Contributions.
-We propose GARLIC, a method for learning geometry-aware representations of high-dimensional spaces via a set of Gaussians optimized for both approximate nearest neighbor retrieval and classification.
-GARLIC learns a probabilistic representation that adapts to local data density and semantic similarity, aligning the shape and placement of Gaussians with the underlying structure of the data.
-In particular, our method adapts to regions with varying intrinsic dimensionality (e.g., curves, surfaces, volumes), enabling each Gaussian to act as an informative geometric primitive and local manifold approximator, rather than a static partitioning cell.
-  • We introduce a geometric representation learning approach inspired by Gaussian Splatting [30], that implicitly partitions the input space, optimized for both retrieval and classification using a unified training objective.
-  • We develop adaptive refinement strategies (split, clone) tailored for high-dimensional spaces, and a localized quantization method for handling density variation, improving search efficiency and robustness w.r.t.
-noise.
-  • We design a novel information-theoretic objective combining Mahalanobis-based divergence, entropy-reducing confidence, and geometry-preserving anchoring, aligning the Gaussian structure with the data manifold.
-  • We benchmark GARLIC for retrieval and classification.
-Our approach exhibits state-of-the-art performance on both tasks w.r.t.
-runtime-recall tradeoff, especially for small candidate sets.
-The source code of our approach will be publicly available.
+Traditional ANN Families.
+ANN methods fall into three main families: 1.
+sketching and compression, which encode vectors into compact codes for fast distance estimation (e.g., product quantization (Jegou et al., 2010), optimized PQ (Ge et al., 2013), iterative quantization (Gong et al., 2012)) 2.
+index-based methods, which pre-organize the dataset to reduce the number of points touched at query time (e.g., IVF (Jegou et al., 2010), PCA-trees (Sproull, 1991), randomized projections and Johnson–Lindenstrauss-based embeddings (Anagnostopoulos et al., 2018) 3.
+graph-based methods, which traverse neighborhood graphs during search (e.g., HNSW (Malkov & Yashunin, 2018), DiskANN (Jayaram Subramanya et al., 2019)).
+Within indices, our work focuses on the sub-family of space-partition indices, which balance memory efficiency with parallelizability and provide a probe-based complexity model compatible with GPU and IO acceleration.
+Data-Independent Partitions.
+Data-independent indices split space according to fixed random rules, ignoring the geometry of the dataset.
+A canonical example is hyperplane LSH, which assigns points based on the sign of random projections and can be queried more flexibly via multi-probing (Indyk & Motwani, 1998, Lv et al., 2007).
+These methods offer theoretical guarantees and fast build times, but their isotropic and geometry-agnostic partitions lead to poor recall on clustered or manifold-structured data.
+GARLIC instead learns anisotropic, density-adaptive cells aligned with the underlying data geometry.
+Data-Dependent Partitions.
+Classical learned indices often rely on isotropic partitions with a fixed global number of cells.
+$k$-Means and its inverted-file variants (IVF) assign points to centroid Voronoi cells (Lloyd, 1982, Jegou et al., 2010), while PCA-trees split recursively along principal components (Sproull, 1991).
+Scalable extensions include mini-batch k-means (Sculley, 2010), BIRCH, which builds a hierarchical clustering tree with compact representations (Zhang et al., 1996), and BLISS, which incrementally refines partitions for large datasets (Gupta et al., 2022).
+These methods are efficient, but their isotropic cells and global resolution fragment dense regions and mix unrelated points in sparse ones.
+Neural LSH takes a different approach by building balanced cuts of the $k$-NN graph and training a classifier to extend them to $\mathbb{R}^{d}$ (Dong et al., 2020).
+While this can outperform k-means in some settings, the emphasis on balance rather than geometry often splits true neighbors in dense areas and merges unrelated points in sparse areas, raising candidate counts.
+Gaussian mixture models (GMMs) capture local covariance through Mahalanobis metrics and soft assignments (Dempster et al., 1977, Banerjee et al., 2005).
+These models demonstrate the benefits of anisotropy, but they maximize likelihood rather than probe efficiency and do not refine capacity locally.
+GARLIC combines the strengths of these directions by learning anisotropic, density-adaptive partitions with local split/clone refinement under an information-theoretic probe-budget objective, explicitly tailored to ANN retrieval.
+The remainder of this work is organized as follows: Section 2 introduces the GARLIC framework, including Gaussian parameterization, the information-theoretic optimization objective, and adaptive refinement strategies.
+Section 3 presents our experimental evaluation on standard Euclidean benchmark datasets, a set of crucial ablation studies, and GARLIC’s limitations.
+Finally, conclusions are drawn in Section 4.
 2 Method
-GARLIC uses a set of $d$-dimensional Gaussians $\mathcal{G}=\{\mathcal{N}(\mu_{i},\Sigma_{i})\}_{i}$, that adapt to the underlying data distribution, provided as a set of points $x\in X\subset\mathbb{R}^{d}$, to represent high-dimensional space for efficient ANNS and classification.
-We choose Gaussians because they capture both local geometry and statistical structure, unlike simpler volumes such as boxes or zonotopes, and express anisotropy, curvature, and density variation in a unified, continuous form.
-The closed form Mahalanobis scoring and differentiability also make them well-suited for end-to-end optimization.
-Starting with data points, we represent the space with the set of Gaussians $\mathcal{G}$ parameterized by means and covariance matrices via Cholesky decomposition (Sec.
-2.1).
-We optimize the coefficients (mean and cholesky) of the Gaussians using objectives that aim to balance coverage, point assignment confidence, and structural consistency (Sec.
-2.2).
-Unlike traditional partitioning methods that enforce or tend towards balanced divisions [14, 19, 6, 31, 1, 41, 39], our approach adaptively refines the set of Gaussians $\mathcal{G}$ through splitting and cloning operations, creating a data-dependent partitioning that accommodates varying densities (Sec.
-2.3).
-To further enhance search efficiency, i.e., narrow down the number of searches, a local hyper-spherical quantization index within each Gaussian bucket (Sec.
-2.4) is built.
-At query time, our method uses Mahalanobis distances for fast bucket selection and prioritized bin search for efficient retrieval (Sec.
-2.5).
-This combination of techniques allows GARLIC to effectively capture the natural clustering structure of high-dimensional data while maintaining search performance, as illustrated in Fig.
-1.
+GARLIC uses a collection of Gaussians $\mathcal{G}=\{\mathcal{N}(\mu_{i},\Sigma_{i})\}_{i}$, whose means and covariances adapt to the underlying data distribution, to partition $X\subset\mathbb{R}^{d}$ into cells for ANN.
+We choose Gaussians because their mean $\mu$ and covariance $\Sigma$ jointly encode geometry and statistics: eigenvectors of $\Sigma$ capture local principal directions, eigenvalues control anisotropy and effective dimensionality, and the Mahalanobis distance $\delta_{M}^{2}(\mathbf{x},g_{i})=(\mathbf{x}-\boldsymbol{\mu}_{i})^{\top}\boldsymbol{\Sigma}_{i}^{-1}(\mathbf{x}-\boldsymbol{\mu}_{i})=\|\mathbf{L}_{i}^{-1}(\mathbf{x}-\boldsymbol{\mu}_{i})\|_{2}^{2}$, is the canonical quadratic form associated with the covariance.
+Level sets of $\delta_{M}$ correspond to $\chi^{2}_{d}$ probability contours, giving a calibrated notion of coverage.
+Parameterizing $\boldsymbol{\Sigma}$ via its Cholesky factor yields closed-form (Section 2.1), differentiable gradients while guaranteeing positive definitenes, which makes Gaussians uniquely well-suited among partitioning primitives for end-to-end optimization, unlike boxes or zonotopes that lack smooth probabilistic distance functions.
+Compared to alternatives such as Gaussian log-likelihood, which adds normalization terms unrelated to retrieval efficiency, or KL divergence, which compares distributions rather than points, Mahalanobis provides a direct, efficient, and anisotropy-aware metric for both training and retrieval.
+We train the Gaussians with information-theoretic objectives that balance coverage, assignment confidence, and structural consistency (Section 2.2), and refine capacity only where needed via split and clone operations, avoiding uniform or balanced partitions that fragment dense regions or merge sparse ones (Dong et al., 2020, Gong et al., 2012, Arthur & Vassilvitskii, 2006, Kumar et al., 2008, Abdullah et al., 2014, Sproull, 1991, McNames, 2001) (Section 2.3).
+At query time, each Gaussian cell is equipped with a local hyper-spherical quantization index to narrow candidate searches (Section 2.4), and Mahalanobis distances are again used for cell selection and prioritized bin search (Section 2.5).
+This approach enables GARLIC to adapt to data geometry while maintaining efficient retrieval, as illustrated in Figure 1.
+[…]
 Figure 1: Overview of GARLIC.
-The method starts with input vector points that are used to initialize a set of Gaussians parameterized by mean $\boldsymbol{\mu}$ and covariance matrices via Cholesky decomposition.
-During optimization, these Gaussians are refined through our novel objectives that balance coverage, assignment confidence, and structural consistency.
-The representation is adaptively improved through split and clone operations creating a data-dependent semantic partitioning that better reflects varying data densities in different regions of the feature space.
+Input vectors are represented by Gaussian cells (means, covariances) optimized with information-theoretic objectives for coverage, confidence, and consistency.
+Cell density is refined adaptively via split/clone operations, and queries use spherical quantization with Mahalanobis distance for cell selection and retrieval.
+Blue arrows indicate transferring information of gradients (back-propagation) or statistics.
 2.1 Gaussian Parameterization
-Each Gaussian $g_{i}$ is defined by its mean $\boldsymbol{\mu_{i}}\in\mathbb{R}^{d}$ and covariance matrix $\boldsymbol{\Sigma}_{i}\in\mathbb{R}^{d\times d}$, parameterized as $\boldsymbol{\Sigma}_{i}=\mathbf{L}_{i}{\mathbf{L}_{i}}^{\top}$, $\mathbf{L}_{i}\in\mathbb{R}^{d\times d}$ being a lower triangular Cholesky factor of matrix $\boldsymbol{\Sigma}_{i}$, ensuring positive semi-definiteness (a requirement for valid Gaussian distributions).
-The means $\boldsymbol{\mu}_{i}$ are initialized using Κ-Μeans++ [7] on a subset of the training set.
-Cholesky factors $\mathbf{L}_{i}$ are initialized as:
-$\mathbf{L}_{i}=\log(\bar{\delta_{\mathcal{E}}})\cdot\mathbf{I}_{d}+\boldsymbol{\epsilon},$
-where $\mathbf{I}_{d}\in\mathbb{R}^{d\times d}$ is the identity matrix, and $\bar{\delta_{\mathcal{E}}}$ is the mean Euclidean distance from each $\boldsymbol{\mu}_{i}$ to its three nearest neighbors.
+Each Gaussian $g_{i}$ is defined by its mean $\boldsymbol{\mu_{i}}\in\mathbb{R}^{d}$ and covariance $\boldsymbol{\Sigma}_{i}\in\mathbb{R}^{d\times d}$, parameterized as $\boldsymbol{\Sigma}_{i}=\mathbf{L}_{i}{\mathbf{L}_{i}}^{\top}$, $\mathbf{L}_{i}\in\mathbb{R}^{d\times d}$ being a lower triangular Cholesky factor of matrix $\boldsymbol{\Sigma}_{i}$, ensuring positive definiteness (a prerequisite for valid Gaussian distributions).
+The means $\boldsymbol{\mu}_{i}$ are initialized using K-Means++ (Arthur & Vassilvitskii, 2007) on a subset of the training set.
+Cholesky factors $\mathbf{L}_{i}$ are initialized as: $\mathbf{L}_{i}=\log(\bar{\delta_{\mathcal{E}}})\cdot\mathbf{I}_{d}+\boldsymbol{\epsilon},$ where $\mathbf{I}_{d}\in\mathbb{R}^{d\times d}$ is the identity matrix, and $\bar{\delta_{\mathcal{E}}}$ is the mean Euclidean distance of each $\boldsymbol{\mu}_{i}$ to its three nearest neighbors.
 This ensures that the initial scale of each Gaussian reflects the local data density.
-The perturbation $\boldsymbol{\epsilon}\in\mathbb{R}^{d\times d}$ is a random lower-triangular matrix, defined below:
-$\boldsymbol{\epsilon}_{jk}=\begin{cases}2\cdot\sigma(r_{jk})-1,&j>k\\ 0,&\text{otherwise}\end{cases}$
-where $r_{jk}\sim\mathcal{U}(0,0.01)$ and $\sigma(\cdot)$ is the sigmoid function.
-This results in diagonal-dominant $\mathbf{L}_{i}$, which stabilizes the optimization process while enabling flexible, anisotropic covariance structures.
-This initialization strategy attempts to align the initial Gaussian scales with the local data density and to ensure sufficient diversity among Gaussians for adaptive refinement in later stages.
+Perturbation $\boldsymbol{\epsilon}\in\mathbb{R}^{d\times d}$ is a random lower-triangular matrix with entries $\epsilon_{jk}=2\sigma(r_{jk})-1$ if $j>k$ and 0 otherwise, where $r_{jk}\sim\mathcal{U}(0,0.01)$ and $\sigma(\cdot)$ is the sigmoid function.
+This construction yields diagonal-dominant $\mathbf{L}_{i}$, stabilizing optimization while allowing anisotropic covariances.
 2.2 Optimization Objective
-We adapt ideas from 3D Gaussian Splatting (3DGS) [29] to high-dimensional nearest neighbor search.
-While 3DGS performs well in 3D scene modeling, its reliance on Structure-from-Motion (SfM) initialization and view-space heuristics breaks down in high dimensions, where geometry is abstract and dimensions encode heterogeneous features.
-Splitting and cloning based on spatial cues become unreliable, and optimization suffers in sparse regions.
-These issues are exacerbated by the curse of dimensionality and clustering, which degrade distance metrics and make balanced partitioning ineffective.
-High-dimensional indexing demands adaptive, data-aware refinement, something 3DGS heuristics cannot provide.
-To handle these issues, inspired by Self-Supervised learning techniques [9, 45], that use information theoretic objectives to increase the capacity of the representation space, we introduce similar information objectives to guide optimization towards a more informative and balanced solution without excessive prior usage.
+High-dimensional indexing requires objectives that remain stable under the curse of dimensionality and robust to heterogeneous feature distributions.
+Classical partitioning heuristics (e.g., balanced splits, uniform clustering) degrade in such regimes: dense regions are fragmented, sparse regions are merged, and distance metrics lose discriminative power Aggarwal et al.
+(2001), Aumüller et al.
+(2020).
+To overcome this, we draw inspiration from self-supervised learning methods such as VICReg Bardes et al.
+(2022) and Barlow Twins Zbontar et al.
+(2021), which employ information-theoretic objectives to increase representational capacity without supervision.
+Analogously, we introduce objectives that guide Gaussians to (i) cover the data distribution, (ii) assign points with high confidence, and (iii) remain anchored to local structure.
+This replaces heuristic spatial rules with principled, differentiable criteria suited to high-dimensional retrieval.
 More specifically, we introduce a divergence-based objective that acts as a reconstruction loss, and regularize it to prevent information explosion (i.e., uncontrolled growth and excessive overlap of Gaussians).
 The divergence loss $\mathcal{L}_{\text{div}}$ is defined as:
-$\mathcal{L}_{\text{div}}=\frac{1}{N}\sum_{\mathbf{x}\in X}\left(\min_{g_{i}\in\mathcal{G}}\delta_{M}(\mathbf{x},g_{i})-\tau\right)^{+},$
-where $\delta_{M}(\mathbf{x},g_{i})=||\mathbf{L}_{i}^{-1}(\mathbf{x}-\boldsymbol{\mu}_{i})||_{2}$ denotes the Mahalanobis distance, $(\cdot)^{+}=\max(0,\cdot)$ and $\tau$ is a standard deviation threshold.
-It penalizes points that fall outside a Gaussian’s coverage radius ($\delta_{M}(\mathbf{x}_{i},g_{i})>\tau$), encouraging Gaussians to expand and cover these points, while points inside ($\delta_{M}(\mathbf{x}_{i},g_{i})\leq\tau$) do not contribute, allowing controlled expansion but preventing infinite growth.
-Still, Gaussians can overlap, leading to redundant information.
+$\mathcal{L}_{\text{div}}=\frac{1}{N}\sum_{\mathbf{x}\in X}\left(\min_{g_{i}\in\mathcal{G}}\delta_{M}(\mathbf{x},g_{i})-\tau\right)^{+},$ (1)
+[…]
+(a) Clone operation.
+(light red) region where $d_{M}(x,g_{i})\leq\tau$; (blue) outer shell $\tau\leq d_{M}(x,g_{i})\leq e\tau$; (gray) $\{x:\ d_{M}(x,g_{i})\geq e\tau\}$; (red) mean of the new gaussian; (green) new covariance matrix.
+(purple - lime) denote increasing data density.
+(b) Grad flow: $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathcal{L}_{\text{div}}$}, $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathcal{L}_{\text{cov}}$}, $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathcal{L}_{\text{anchor}}$}
+Figure 2: Gaussian refinement and associated loss gradients.
+Still, Gaussians can overlap, leading to redundant information and performance loss.
 To surpass this issue, and mitigate fuzzy assignments of points to Gaussians, we introduce a covariance-based regularization, which encourages each Gaussian to dominate its assigned points.
-Specifically, given a point $\mathbf{x}$, we define its coverage set, i.e., the set of Gaussians that satisfy the coverage radius constraint, as $\mathcal{M}(\mathbf{x})=\{g_{i}\in\mathcal{G}\mid\delta_{M}(\mathbf{x},g_{i})\leq\tau\}.$
-Then, we compute the normalized soft-assignment probabilities based on Euclidean distances ($\delta_{\mathcal{E}}$) as $p_{i}(\mathbf{x})=e^{-\delta_{\mathcal{E}}(\mathbf{x},\boldsymbol{\mu}_{i})}/\sum_{g_{j}\in\mathcal{M}(\mathbf{x})}e^{-\delta_{\mathcal{E}}(\mathbf{x},\boldsymbol{\mu}_{j})}+\epsilon,\ \forall g_{i}\in\mathcal{M}(\mathbf{x})$.
+Specifically, given a point $\mathbf{x}$, we define its coverage set, i.e., the set of Gaussians that satisfy the coverage radius constraint, as $\mathcal{M}(\mathbf{x})=\{g_{i}\in\mathcal{G}\mid\delta_{M}(\mathbf{x},g_{i})\leq\tau\}.$ Then, we compute the normalized soft-assignment probabilities based on Euclidean distances ($\delta_{\mathcal{E}}$) as $p_{i}(\mathbf{x})=e^{-\delta_{\mathcal{E}}(\mathbf{x},\boldsymbol{\mu}_{i})}/\sum_{g_{j}\in\mathcal{M}(\mathbf{x})}e^{-\delta_{\mathcal{E}}(\mathbf{x},\boldsymbol{\mu}_{j})}+\epsilon,\ \forall g_{i}\in\mathcal{M}(\mathbf{x})$.
 The covariance loss $\mathcal{L}_{\text{cov}}$ is defined as:
-$\mathcal{L}_{\text{cov}}=1-\frac{1}{N}\sum_{\mathbf{x}\in X}\max_{g_{i}\in\mathcal{M}(\mathbf{x})}p_{i}(\mathbf{x})$
+$\mathcal{L}_{\text{cov}}=1-\frac{1}{N}\sum_{\mathbf{x}\in X}\max_{g_{i}\in\mathcal{M}(\mathbf{x})}p_{i}(\mathbf{x})$ (2)
 This loss encourages highly confident (low-entropy) assignments, thereby reducing ambiguity and stabilizing optimization.
-To prevent excessive expansion of Gaussians caused by the divergence objective and ensure that each Gaussian accurately aligns with its assigned points, we introduce the anchor loss $\mathcal{L}_{\text{anchor}}$:
-$\mathcal{L}_{\text{anchor}}=\frac{1}{d|G|}\sum_{g_{i}\in G}\left(||\boldsymbol{\mu}_{i}-\hat{\boldsymbol{\mu}}_{i}||_{2}^{2}+\alpha||\mathbf{L}_{i}{\mathbf{L}_{i}}^{\top}-\hat{\boldsymbol{\Sigma}}_{i}||_{F}^{2}\right),$
-where $\hat{\boldsymbol{\mu}}_{i}$ and $\hat{\boldsymbol{\Sigma}}_{i}=\mathrm{Cov}(x\in B_{g_{i}})$ are the empirical mean and covariance of points assigned to Gaussian $g_{i}$, and $\alpha$ is a hyperparameter balancing position and shape.
-This loss constrains Gaussians by anchoring them closely to their local data distributions, counteracting uncontrolled growth from other loss terms and maintaining geometric alignment and interpretability.
+To further prevent excessive expansion of Gaussians caused by the divergence objective and ensure that each Gaussian aligns with its assigned points, we introduce the anchor loss $\mathcal{L}_{\text{anchor}}$:
+[…]
+where $\hat{\boldsymbol{\mu}}_{i}$ and $\hat{\boldsymbol{\Sigma}}_{i}=\mathrm{Cov}(x\in\mathcal{B}_{g_{i}})$ are the empirical mean and covariance of points assigned to Gaussian $g_{i}$, and $\alpha$ is a hyperparameter balancing position and shape.
+This loss constrains Gaussians by anchoring them closely to their local data distributions, restraining uncontrolled growth from other loss terms and maintaining geometric alignment.
 Finally, our loss function is defined as:
@@
 where $\lambda_{\text{div}},\lambda_{\text{cov}}\text{ and }\lambda_{\text{anchor}}$ are hyperparameters balancing the importance of each term.
-(a) Clone operation.
-(light red) region where $d_{M}(x,g_{i})\leq\tau$; (blue) outer shell $\tau\leq d_{M}(x,g_{i})\leq e\tau$; (gray) $\{x:\ d_{M}(x,g_{i})\geq e\tau\}$; (red) mean of the new gaussian; (green) new covariance matrix.
-(purple - lime) denote increasing data density.
-(b) Grad flow: ${\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\mathcal{L}_{\text{div}}}$, ${\color[rgb]{0.72,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.72,0,0}\pgfsys@color@cmyk@stroke{0}{0.89}{0.94}{0.28}\pgfsys@color@cmyk@fill{0}{0.89}{0.94}{0.28}\mathcal{L}_{\text{cov}}}$, ${\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}\mathcal{L}_{\text{anchor}}}$
-Figure 2: Gaussian refinement and associated loss gradients.
 2.3 Adaptive Refinement
@@
 We employ an adaptive refinement strategy that adjusts the Gaussian set based on local point density.
-Prior approaches rely on positional gradients to guide refinement [29], but these become sparse and unreliable in high dimensions.
-In our case, we also require search efficiency, so we refine Gaussians whose bucket cardinality exceeds a threshold: $|\mathcal{B}_{g}|>\gamma\cdot|X|$.
-Splitting is performed by applying DBSCAN [17] to detect two dense clusters within the bucket.
-If unsuccessful, we fall back to K-Means to produce new means $\boldsymbol{\mu}_{1},\boldsymbol{\mu}_{2}$, while covariances are scaled down as $\mathbf{L}_{1}=\mathbf{L}_{2}=\alpha\cdot\mathbf{L}$, with $\alpha<1$.
+Prior approaches rely on positional gradients to guide refinement (Kerbl et al., 2023), but these become sparse and unreliable in high dimensions.
+In our case, we also require search efficiency, so we refine Gaussians whose cell cardinality exceeds a threshold: $|\mathcal{B}_{g}|>\gamma\cdot|X|$.
+Splitting is performed by applying clustering (Arthur & Vassilvitskii, 2007) to produce new means $\boldsymbol{\mu}_{1},\boldsymbol{\mu}_{2}$, while covariances are scaled down as $\mathbf{L}_{1}=\mathbf{L}_{2}=\alpha\cdot\mathbf{L}$, with $\alpha<1$.
 For cloning, we target Gaussians with a high ratio of outside points to interior ones $\mathcal{B}_{g_{i}}^{in}=\{\mathbf{x}\in X:\delta_{M}(\mathbf{x},g_{i})\leq\tau\}$.
-Instead of using all nearest outside points, we focus on a boundary region, and select $k=\rho\cdot|\mathcal{B}_{g}^{out}|$ points, where $\rho\in(0,1)$ is a sampling ratio and $\mathcal{B}_{g_{i}}^{out}=\{\mathbf{x}\in X:\tau<\delta_{M}(\mathbf{x},g_{i})<e\tau\text{ and }g_{i}=\arg\min_{g_{j}\in G}\delta_{M}(\mathbf{x},g_{j})\}$ with $e>1$ is the set of all boundary points.
+Instead of using all nearest outside points, we focus on a boundary region, and select $k=\rho\cdot|\mathcal{B}_{g}^{out}|$ points, where $\rho\in(0,1)$ is a sampling ratio and $\mathcal{B}_{g_{i}}^{out}=\{\mathbf{x}\in X:\tau<\delta_{M}(\mathbf{x},g_{i})<e\tau\text{ and }g_{i}=\arg\min_{g_{j}\in\mathcal{G}}\delta_{M}(\mathbf{x},g_{j})\}$ with $e>1$ is the set of all boundary points.
 The sampled subset $S=\{\mathbf{x}_{1},\mathbf{x}_{2},...,\mathbf{x}_{k}\}\subset\mathcal{B}_{g}^{out}$ is chosen randomly without replacement.
-This sampling approach serves several purposes, reducing computational cost, as density estimation in high dimensions is expensive, providing statistical robustness by focusing on representative points and helping to avoid outliers that might exist in the boundary region as Fig.
-2 showcases.
-The candidate Gaussians are picked with regards to the ration of boundary to interior points: $|\mathcal{B}_{g}^{out}|/|\mathcal{B}_{g}^{in}|>\beta$.
-The center of the new Gaussian is placed at the point with the highest local density in the boundary region, such that for a point $\mathbf{p}$ in the boundary region, we compute its local density as:
-$\rho(\mathbf{p})=\frac{1}{\frac{1}{k}\sum_{j=1}^{k}\delta_{\mathcal{E}}(\mathbf{p},\text{NN}_{j}(\mathbf{p}))},$
-here $\text{NN}_{j}(\mathbf{p})$ is the $j$-th nearest neighbor of $\mathbf{p}$ and $\delta_{\mathcal{E}}(\cdot,\cdot)$ is the Euclidean distance.
-We select the point with the highest density: $\mathbf{p}^{*}=\arg\max_{\mathbf{p}}\rho(\mathbf{p})$.
+This sampling approach serves several purposes, reducing computational cost, as density estimation in high dimensions is expensive, providing statistical robustness by focusing on representative points and helping to avoid outliers that might exist in the boundary region as Figure 2 showcases.
+The candidate Gaussians are picked with regard to the ratio of boundary to interior points: $|\mathcal{B}_{g}^{out}|/|\mathcal{B}_{g}^{in}|>\beta$.
+The center of the new Gaussian is placed at the point with the highest local density in the boundary region, such that for a point $\mathbf{p}$ in the boundary region, we compute its local density as the inverse mean Euclidean distance from the $3$-NN and select the point with the highest density: $\mathbf{p}^{*}=\arg\max_{\mathbf{p}}\rho(\mathbf{p})$.
 The covariance matrix is then cloned, with $\mathbf{L}_{new}=\mathbf{L}$.
+In our method, splitting is prioritized over cloning: when a Gaussian grows too large, we first partition it to reduce cell cardinality and maintain search efficiency.
+Cloning is applied only if splitting is not triggered, serving to refine coverage near dense boundary regions without inflating candidate counts unnecessarily.
 2.4 Quantization
-After optimizing our Gaussians, we assign points that fall outside the coverage radius to the nearest Gaussian such that our final buckets are:
-$B_{g_{i}}=\{\mathbf{x}\in X\mid\delta_{M}(\mathbf{x},g_{i})\leq\tau\}\cup\{\mathbf{x}\in X\mid\delta_{M}(\mathbf{x},g_{i})>\tau\text{ and }g_{i}=\arg\min_{g_{j}\in G}\delta_{M}(\mathbf{x},g_{j})\}$
-Algorithm 1 Build Index
-1:Input: Dataset $\mathcal{D}$ of $d$-dimensional points
-2:Initialize Gaussians on $\mathcal{D}$
-3:$\mathcal{G}\leftarrow$ Obtain $K$ Gaussians $g_{i}=\mathcal{N}(\mu_{i},L_{i})$ (Sec. 2.22.3)
-4:for each point $p$ do
-5:    Assign $p$ to all Gaussians where $d_{M}(p,g_{i})<\tau$, $\forall i\in[K]$
-6:    if $p$ not assigned to any Gaussian then
-7:         Assign $p$ to nearest Gaussian w.r.t. $d_{M}$
-8:    end if
-9:end for
-10:for each Gaussian $g_{i}\in\mathcal{G}$ do
-11:    $\mathcal{Q}_{i}\leftarrow$ Quantize indexed data of $g_{i}$ (Sec. 2.4)
-12:end for
-13:return $\{(g_{i},\mathcal{Q}_{i})\}_{i\in[K]}$
-While our Gaussian-based buckets provide an effective organization of the data space, exhaustive search within buckets becomes prohibitively expensive as dimensionality increases.
-For high-dimensional nearest neighbor search, we must avoid calculating distances to all points within candidate buckets.
-Standard approximate search methods, however, often do not account for the unique structure provided by our Gaussian-based partitioning.
-We introduce a quantization scheme that leverages the inherent properties of our Gaussian model.
-The key insight is that once we have assigned points to appropriate Gaussian buckets, we can apply efficient Euclidean-based quantization within each bucket’s local coordinate system, reducing the number of distance computations required at query time.
-Algorithm 2 Query Index
-1:Input: Index structure set $\{(g_{i},\mathcal{Q}_{i})\}_{i\in[K]}$, query $q$
-2:$g_{i}\leftarrow\arg\min_{i:g_{i}\in\mathcal{G}}d_{M}(q,g_{i})$
-3:$B\leftarrow$ Traverse subset of bins from $\mathcal{Q}_{i}$ (Sec. 2.4)
-4:Measure $L_{2}$ distance to candidates in $B$
-5:Terminate if condition (Sec. 2.5)
-6:return nearest neighbor of $q$ in $B$
-For each bucket $B_{g}$, we apply PCA to reduce dimensionality while preserving the local structure $\mathbf{P}_{g}=\text{PCA}(\{\mathbf{x}-\bar{\mathbf{x}}_{g}\mid\mathbf{x}\in B_{g}\},r)$ where $\bar{\mathbf{x}}_{g}$ is the mean of points in bucket $B_{g}$, $r$ is the reduced dimensionality (typically $r\ll d$, constant in practice), and $\mathbf{P}_{g}\in\mathbb{R}^{d\times r}$ contains the principal components.
-Each point $\mathbf{x}\in B_{g}$ is projected into this lower-dimensional space $\mathbf{x}^{r}=\mathbf{P}_{g}^{T}(\mathbf{x}-\bar{\mathbf{x}}_{g})$, after which its transformed to hyperspherical coordinates $\mathbf{s}=\text{cart2sph}(\mathbf{x}^{r})$, where the transformation for an $r$-dimensional point yields coordinates $\mathbf{s}=(s_{1},s_{2},\ldots,s_{r})$.
-The hyperspherical space is then partitioned into bins defined by $n_{r}$ radial divisions and $n_{a}$ angular divisions per dimension: $\mathcal{B}_{g}=\{B_{i,\mathbf{j}}\mid i\in\{1,\ldots,n_{r}\},\mathbf{j}\in\{1,\ldots,n_{a}\}^{r-1}\}$ with boundaries $r_{i}=r_{\min}+(r_{\max}-r_{\min})\cdot i/n_{r},\ i\in\{0,\ldots,n_{r}\}$ and $\theta_{j,k}=\theta_{\min,k}+(\theta_{\max,k}-\theta_{\min,k})\cdot j/n_{a},\ j\in\{0,\ldots,n_{a}\},k\in\{1,\ldots,r-1\}$.
-Alg.
-1 illustrates the construction process of the index structure.
+After optimizing our Gaussians, we assign points that fall outside the coverage radius to the nearest Gaussian such that our final cells are: $\mathcal{B}_{g_{i}}=\{\mathbf{x}\in X\mid\delta_{M}(\mathbf{x},g_{i})\leq\tau\}\cup\{\mathbf{x}\in X\mid\delta_{M}(\mathbf{x},g_{i})>\tau\text{ and }g_{i}=\arg\min_{g_{j}\in G}\delta_{M}(\mathbf{x},g_{j})\}$.
+While Gaussian cells provide an effective partition of the space, brute-force search within them is still prohibitive in high dimensions.
+Standard ANN methods usually quantize globally, ignoring the anisotropy and local geometry already captured by our Gaussians.
+We instead introduce a localized quantization scheme: once points are assigned to cells, each cell is treated in its own coordinate system, where Euclidean structure is better aligned with the underlying data.
+Quantizing in this local basis reduces the number of distance computations per query, while preserving the geometry captured by Mahalanobis distance.
+For each cell $\mathcal{B}_{g}$, we apply PCA to reduce dimensionality while preserving the local structure $\mathbf{P}_{g}=\text{PCA}(\{\mathbf{x}-\bar{\mathbf{x}}_{g}\mid\mathbf{x}\in\mathcal{B}_{g}\},r)$ where $\bar{\mathbf{x}}_{g}$ is the mean of points in cell $\mathcal{B}_{g}$, $r$ is the reduced dimensionality (typically $r\ll d$, constant in practice), and $\mathbf{P}_{g}\in\mathbb{R}^{d\times r}$ contains the top-$r$ principal components.
+This step both lowers computational cost and aligns the local coordinate system with the main variance directions of the data.
+Each point $\mathbf{x}\in\mathcal{B}_{g}$ is then projected into this subspace $\mathbf{x}^{r}=\mathbf{P}_{g}^{\top}(\mathbf{x}-\bar{\mathbf{x}}_{g})$, which embeds the data in $\mathbb{R}^{r}$ while preserving Euclidean structure up to the discarded components.
+Finally, we convert $\mathbf{x}^{r}$ into hyperspherical coordinates $\mathbf{s}=\text{cart2sph}(\mathbf{x}^{r})=(s_{1},s_{2},\ldots,s_{r})$, where $s_{1}$ denotes the radial component $\|\mathbf{x}^{r}\|_{2}$ and $s_{2},\dots,s_{r}$ are angular coordinates.
+This reparameterization retains the Euclidean metric but enables partitioning along radial and angular directions.
+The hyperspherical space is partitioned into bins $\mathcal{Q}_{g}=\{B_{i,\mathbf{j}}\mid i\in\{1,\ldots,n_{r}\},\;\mathbf{j}\in\{1,\ldots,n_{a}\}^{\,r-1}\}$, with radial boundaries $r_{i}=r_{\min}+(r_{\max}-r_{\min})\tfrac{i}{n_{r}},\;i=0,\ldots,n_{r}$, and angular boundaries $\theta_{j,k}=\theta_{\min,k}+(\theta_{\max,k}-\theta_{\min,k})\tfrac{j}{n_{a}},\;j=0,\ldots,n_{a},\;k=1,\ldots,r-1$.
+The dominant cost of index construction comes from full Mahalanobis-based assignments during optimization, requiring $\mathcal{O}(|X|\cdot K\cdot d^{2})$ time per iteration, while initialization, refinement, and PCA quantization add only lower-order terms (see Appendix A.3).
 2.5 Inference
-At query time, for a query point $\mathbf{q}$, we employ a multi-stage approach to efficiently retrieve approximate nearest neighbors.
-First, we identify the $k$ nearest Gaussians using the Mahalanobis distance.
-For each selected Gaussian bucket $g$, we project the query into the bucket’s local PCA space, $\mathbf{q}^{r}_{g}=\mathbf{P}_{g}^{T}(\mathbf{q}-\bar{\mathbf{x}}_{g})$, and convert it to hyperspherical coordinates $\mathbf{s}_{g}=\text{cart2sph}(\mathbf{q}^{r}_{g})$.
-We then compute the distances to all bins in the bucket and sort them in ascending order.
-For each bin $B_{i,\mathbf{j}}$ with boundaries defined by $r$-intervals $[r_{i},r_{i+1}]$ and angular intervals $[\theta_{j,k},\theta_{j+1,k}]$, we compute the distance $d(\mathbf{s}_{g},B_{i,\mathbf{j}})=\min_{\mathbf{y}\in B_{i,\mathbf{j}}}\|\mathbf{s}_{g}-\mathbf{y}\|_{2}.$
-This minimization problem is efficiently solved by first checking if $\mathbf{s}_{g}$ is inside the bin.
-If it is outside, we solve the constrained optimization problem by initializing a point $\mathbf{y}^{0}$ with its radial component clipped to $[r_{\min},r_{\max}]$ and angular components set to the midpoints of their respective intervals.
-We then minimize $\|\mathbf{s}_{g}-\mathbf{y}\|_{2}$ subject to the bin boundary constraints using the L-BFGS-B [10] algorithm, which respects the bounds $r_{\min}\leq y_{1}\leq r_{\max}$ and $\theta_{\min,j}\leq y_{j+1}\leq\theta_{\max,j}$ for the radial and angular components respectively.
-We then search for points in bins sequentially according to their sorted distances, computing exact Euclidean distances between the query and the points in each bin.
-The search terminates under the condition that we have examined a predefined percentage of bins (typically 30%).
-Given a query, Alg.
-2 showcases the procedure of retrieving the nearest neighbor using our index structure.
+Given a query $\mathbf{q}$, we first select the top-$k_{G}$ Gaussians by Mahalanobis score $\delta_{M}(\mathbf{q},g)$.
+For each selected Gaussian $g$, we project $\mathbf{q}$ to its local PCA space $\mathbf{q}^{r}_{g}=\mathbf{P}_{g}^{\top}(\mathbf{q}-\bar{\mathbf{x}}_{g})$ and convert to hyperspherical coordinates $\mathbf{s}_{g}=\text{cart2sph}(\mathbf{q}^{r}_{g})$.
+We prioritize bins by a query-to-bin distance computed in the reduced Euclidean space.
+Each bin $B_{i,\mathbf{j}}$ is defined by spherical bounds $r\in[r_{i},r_{i+1}]$ and $\theta_{k}\in[\theta_{j_{k},k},\theta_{j_{k}+1,k}]$ for $k=1,\dots,r-1$.
+We define:
+[…]
+If $\mathbf{s}_{g}$ lies inside $B_{i,\mathbf{j}}$, then $d(\mathbf{q}^{r}_{g},B_{i,\mathbf{j}})=0$; otherwise we solve the bound-constrained problem (Byrd et al., 1995) initialized at $\phi_{1}=\text{clip}(\|\mathbf{q}^{r}_{g}\|_{2},[r_{i},r_{i+1}])$ and $\phi_{k+1}=0.5(\theta_{j_{k},k}+\theta_{j_{k}+1,k})$.
+Bins in $\mathcal{Q}_{g}$ are sorted by $d(\mathbf{q}^{r}_{g},\cdot)$ ascending and scanned until a bin budget $\rho\in(0,1]$ of bins per cell is exhausted (typically $\rho=0.3$).
+Within each visited bin we compute exact Euclidean distances in the original space between $\mathbf{q}$ and all points indexed in that bin, aggregating candidates across the $k_{G}$ selected Gaussians.
+The final result is obtained by selecting the overall nearest neighbors among the accumulated candidates, with the dominant inference cost coming from distance computations (see Appendix A.3).
+OursK-MeansN-LSHITQCP-LSH
+HP-LSHPCA TreeFaiss-IVFFaiss-IVFPQFS
+BIRCHGMMOurs 5%Ours 2.5%Ours 1%
+(c) Efficiency (Sift)
+(d) SIFT1M
+(e) MNIST
+(f) Fashion-MNIST
+[…]
 3 Experiments
 Datasets.
-We evaluate our method on three benchmark datasets: SIFT1M [37] (128-dimensional image descriptors with one million points), MNIST [32] and Fashion-MNIST [44] (784-dimensional vectors from 28×28 grayscale images).
-These datasets provide diverse characteristics in terms of extrinsic and intrinsic dimensionality, scale, and data distribution, offering a comprehensive evaluation framework in Euclidean space.
-Additional information about the datasets and their licenses are provided in the Supplementary Material (SM).
-OursK-MeansN-LSHITQCP-LSH
+We evaluate our method on three benchmark datasets: SIFT1M (Lowe, 2004) (128-dimensional image descriptors with one million points), MNIST (LeCun et al., 1998) and Fashion-MNIST (Xiao et al., 2017) (784-dimensional vectors from $28\times 28$ grayscale images) following the standard setup in ANN-Benchmarks (Aumüller et al., 2020).
+Further information about the datasets can be found in the Appendix.
+Evaluation.
+We evaluate GARLIC on the approximate nearest neighbor task, reporting Recall@1 and as function of the number of distance computations while varying the probe budget.
+Recall@1 measures the accuracy of the very first retrieved neighbor, reflecting ranking performance, and is strict.
+In Appendix A.1, we showcase more related metrics such as $\epsilon$-Recall, empricial c-approximation factors and mean relative error.
+We use distance computations, rather than wall-clock latency (QPS), as the efficiency axis.
+Distance computations capture the algorithmic effort of a search procedure independently of hardware and implementation details (different CPUs/GPUs, BLAS kernels, batching, I/O) (Aumüller et al., 2020, Peng et al., 2023).
+This metric directly reflects the goal of space-partitioning methods, which is to minimize the number of distances required to achieve a target recall.
+(a) Generalization
+(b) Batch size
+(c) $|\mathcal{G}|$ impact
+(d) Initialization
+(e) Split threshold $\gamma$
+(f) Clone threshold $e$
+[…]
+Baselines and Comparisons.
+We evaluate GARLIC against representative approximate nearest neighbor methods such as $k$-Means++ (Arthur & Vassilvitskii, 2006), BIRCH (Zhang et al., 1996), Gaussian Mixture Models (GMM) (Dempster et al., 1977, Banerjee et al., 2005), Neural-LSH (N-LSH) (Dong et al., 2020), Hyperplane LSH (HP-LSH) (Indyk & Motwani, 1998), Cross-polytope LSH (CP-LSH) (Andoni et al., 2015), ITQ (Gong et al., 2012), PCA Tree (Kumar et al., 2008, Abdullah et al., 2014, Sproull, 1991, McNames, 2001), Faiss-IVF and Faiss-IVFPQFS (Douze et al., 2024).
+All baselines were run under identical conditions, using the standardized ANN-Benchmarks datasets and query splits, with hyperparameters optimized according to their respective publications and usage.
+Results and Analysis.
+We present performance curves across datasets, reporting Recall@1 as a function of the number of retrieved candidates.
+As shown in Figure 3, our method performrs efficiently with respect to other baselines.
+GARLIC generally achieves preferable recall–efficiency trade-offs than traditional space partitioning methods, including inverted-file indices (Faiss-IVF, IVFPQFS), hashing approaches (HP-LSH, CP-LSH, Neural-LSH), and tree-based methods (PCA-Tree, BIRCH).
+Figure 3(a) evaluates efficiency through Recall@1-per-probe versus build time.
+All GARLIC variants occupy the upper-left region, indicating favorable trade-offs: lightweight models trained with $1\%$ of the data already surpass strong baselines, while larger-capacity variants (e.g., $10\%$ training) further improve recall without excessive cost.
+In contrast, competing methods cluster in lower-efficiency or higher-cost regions, reflecting less balanced trade-offs between indexing overhead and search quality.
+Ablation study.
+We conduct ablation studies to isolate the effect of key design choices and components.
+Unless stated otherwise, experiments are on SIFT-1M with a training subset of $5\%$ of $|X|$, and Recall@1 is measured against distance computations.
+We investigate the robustness and data efficiency of our method through controlled downsampling of the training set.
+As shown in Figure 4(4(a)), our method consistently maintains strong performance across varying training sizes.
+Notably, even with only $1\%$ of the training data, our model outperforms the Faiss-IVF variants of up to $10\%$ training size in terms of Recall@1 versus the number of retrieved candidates.
+This trend remains consistent across higher percentages, demonstrating that our approach learns a compact yet highly effective representation of the data distribution.
+Furthermore, Figure 4(4(b)) highlights the batch-size effect (here using a fixed training subset of $100k$ points instead of $50k$), where larger batches yield consistently better performance.
+Finally, Figure 4(4(c)) shows that more initial Gaussians produce better results at equal iterations, reflecting faster convergence.
+This happens because adaptive density control targets spatial placement rather than Gaussian count.
+Table 1: Effect of loss terms and adaptive refinement average Recall@1 / Candidates ($\times 10^{5}$). Higher is better.
+Configuration Performance $\uparrow$
+Loss terms
+   w/ all 16.20
+   $\mathcal{L}_{div}+\mathcal{L}_{cov}$ 15.38
+   $\mathcal{L}_{div}+\mathcal{L}_{anchor}$ 10.64
+   $\mathcal{L}_{div}$ 19.78
+[…]
+   w/ both 16.20
+   w/ split 14.22
+   w/ clone 15.45
+   w/o any 14.63
+We next evaluate core design hyperparameters.
+In Figure 4(4(d)), we investigate initialization strategies for Gaussian centers, revealing that K-Means initialization consistently outperforms alternatives.
+Although farthest point sampling (FPS) achieves comparable recall, it requires approximately 50% more probes, and catching up to K-Means requires more training time.
+In Figure 4(4(e)), the split threshold $\gamma$, controlling the density condition for Gaussian subdivision, is examined.
+The optimal values are $\gamma=0.005$ and $\gamma=0.01$, though $\gamma=0.005$ requires more time due to more frequent splits.
+As for the clone threshold $e$, depicted in Figure 4(4(f)), which determines the outer shell for cloning, it remains stable across different values, with $e=2.2$ being preferable in situations requiring high recall.
+Table 1 analyzes the impact of loss terms, adaptive refinement and covariance configurations.
+Split and clone are pivotal for performance, while the covariance term ($\mathcal{L}_{\text{cov}}$) significantly boosts retrieval.
+The anchor term ($\mathcal{L}_{\text{anchor}}$) has minor effect on raw performance but is essential for making Gaussians geometrically informative.
+Additional ablations and results are provided in the Appendix.
+Limitations and Future Work.
+GARLIC employs Mahalanobis distance and Gaussian primitives, which assume a Euclidean metric and are therefore not directly compatible with angular similarity.
+This may be addressed by adopting angular counterparts, such as von Mises–Fisher distributions, which we leave for future work.
+The trade-off between recall and latency can be further enhanced by augmenting the number of Gaussians and organizing them within tree structures (e.g., KD-tree or Ball-tree over Gaussian means), reducing query cost.
+The spatial complexity scales quadratically with dimensionality ($\mathcal{O}(Kd^{2})$) due to full anisotropic covariances; this challenge can be mitigated through low-rank approximations or quantization of Cholesky factors.
+Finally, while the method shows some sensitivity to initialization, this is alleviated by the information-theoretic objectives and the progressive refinement via split and cloning.
+Beyond these limitations, GARLIC is naturally incremental: each Gaussian cell maintains sufficient statistics (mean, covariance, cardinality), allowing new data to be integrated through online updates.
+Combined with local split/clone refinement, this enables streaming and online learning scenarios without rebuilding the entire index.
+We leave a systematic evaluation of this capability to future work.
+4 Conclusions
+We introduced GARLIC, a geometric structure that learns the underlying distribution for both approximate nearest neighbor search and classification.
+By combining information-theoretic objectives with adaptive refinement (split and clone), and representing the space via anisotropic Gaussians, our method achieves competitive performance in Euclidean approximate nearest neighbor, particularly in low-probe regimes.
+Experiments demonstrate competitive recall-efficiency tradeoffs and robustness under severe data reduction, highlighting its generalization capabilities.
+Appendix A Appendix
+This appendix provides additional details and experiments as mentioned in the main paper.
+Section A.1 provides additional results and ablation studies to the ones presented in Section 3 of the main paper.
+Section A.1.1 contains visualization of diagnostics and statistics regarding our proposed method, to provide a comprehensive understanding of its behavior.
+Section A.2 contains additional technical details regarding the datasets used and the experimental setup, as well as the training pipeline.
+Section A.3 discusses and provides detailed computational time and space complexity for the build and query procedures of GARLIC.
+A.1 Further Results
+[…]
+OursK-MeansBIRCHCP-LSH
 HP-LSHPCA TreeFaiss-IVFFaiss-IVFPQFS
-Ours 10%Ours 5%Ours 2.5%Ours 1%
-(a) Efficiency
-(b) MNIST
-(c) Fashion-MNIST
-(d) SIFT1M
-(e) SIFT1M
-(f) MNIST
-(g) Fashion-MNIST
 […]
-Evaluation.
-We evaluate our method on retrieval and classification tasks across the aforementioned datasets.
-For retrieval, we report Recall10@10 and Recall@1 versus the number of retrieved candidates (probes), reflecting the trade-off between accuracy and efficiency.
-Recall10@10 measures how many of the true top-10 neighbors are in the retrieved set, while Recall@1 checks if the top-ranked result is the true nearest neighbor.
-The candidate count controls query time and serves as the main efficiency metric.
-For MNIST and Fashion-MNIST, we assess classification via majority voting over retrieved candidates: $\hat{y}(q)=\arg\max_{c}\sum_{x\in C_{q}}\mathbb{I}[\text{label}(x)=c]$.
-This evaluates how well the partitioning preserves semantic similarity.
-Baselines and Comparisons.
-We evaluate our proposed method against several state-of-the-art approximate nearest neighbor search algorithms such as K-Means [6], Neural-LSH (N-LSH) [14], Hyperplane LSH (HP-LSH) [25, 5], Cross-polytope LSH (CP-LSH) [5], ITQ [19], PCA Tree [31, 1, 41, 39], and Faiss-IVF and Faiss-IVFPQFS [15].
-For classification tasks, we additionally compare against standard $k$-Nearest Neighbors ($k$-NN) [12, 42], which performs exhaustive search for exact neighbors, and $k$-K-Nearest Neighbors ($kK$-NN) [33], which adjusts the neighborhood size based on local curvature estimates.
-Each method was evaluated under identical conditions, with hyperparameters optimized according to their original publications to ensure fair and meaningful comparisons across both nearest neighbor retrieval and classification performance.
-Table 1: Classification accuracy via majority voting.
-Ours-1: $\arg\min_{g}\delta_{M}(q,g)$; Ours-2: $\delta_{M}(q,g)\leq\tau$; Ours-3: top-$k$ Gaussians.
-We highlight the best, second-best and third-best results.
-Dataset
-Method MNIST $\uparrow$ Fashion-MNIST $\uparrow$
-k-NN 0.7576 0.7133
-kK-NN 0.7609 0.7796
-K-Means 0.9118 0.7589
-PCA Tree 0.7750 0.7427
-ITQ 0.7156 0.6459
-HP-LSH 0.5483 0.6643
-CP-LSH 0.7801 0.7128
-N-LSH 0.4142 0.4307
+Refer to captionMNIST
+Refer to captionFMNIST
+[…]
+Refer to captionPrecision @ 10
+[…]
+We evaluate retrieval quality using both accuracy- and distortion-based criteria, as shown in Figures 5 and 6.
+Beyond Recall@$1$, we report Precision@10 (P@10), capturing the fraction of retrieved points among the top-10 that are true neighbors.
+To assess approximation tightness for the nearest neighbor, we measure the distance ratio between the returned neighbor and the exact nearest neighbor.
+From this we derive: (i) $\varepsilon$-Recall, the fraction of queries where the retrieved distance is within a factor $(1+\varepsilon)$ of the ground-truth (with $\varepsilon\in\{0.01,0.10\}$); (ii) the percentiles $r95$ and $r99$, which correspond to empirical $c$-approximation factors at the 95th and 99th percentiles; and (iii) the mean relative error, $RE_{\text{mean}}$, summarizing the average stretch beyond the true nearest-neighbor distance.
+These metrics complement recall by quantifying how close the returned distances are to the exact nearest neighbor, and follow standard practice in ANN-Benchmarks.
+OursK-MeansBIRCHCP-LSH
+HP-LSHPCA TreeFaiss-IVFFaiss-IVFPQFS
+Refer to captionr95SIFT
+Refer to captionMNIST
+Refer to captionFMNIST
+Refer to captionr99
 […]
 […]
-[…]
-[…]
-Ours-$3$ 0.9175 0.7879
-Results and Analysis.
-We present performance curves across datasets, reporting Recall10@10 and Recall@1 as a function of the number of retrieved candidates.
-As shown in Fig.
-3, our method consistently outperforms state-of-the-art baselines across both metrics.
-Notably, it exhibits a significant advantage in the low-probe regime: for example, on SIFT1M, our method achieves Recall@1 of $\sim 69\%$ using $\sim 3400$ candidates, compared to 61% for Faiss-IVF and 55% for Faiss-IVFPQFS using similar candidate counts.
-This efficiency gap is most evident below the 80% recall threshold, where traditional methods struggle to match the accuracy without a sharp rise in probe cost.
-Fig.
-3(a) highlights the efficiency trade-off between training time and retrieval quality, measured via Recall@1-per-probe.
-All our variants lie in the top-left region of the plot, offering options ranging from ultra-fast lightweight models (e.g., Ours 1%) to more accurate, higher-capacity variants (e.g., Ours 10%).
-In contrast, competing methods tend to occupy lower-efficiency or higher-cost zones, demonstrating a less favorable balance between indexing overhead and retrieval quality.
-(a) Generalization
-(b) Time-recall@1
-(c) Quantization
-(d) $|\mathcal{G}|$ impact
-Figure 4: (a) Generalization with limited training data. (b) Time-recall@1 tradeoff. (c) Quantization: efficient search via partial bucket scanning. (d) Impact of initial number of Gaussians.
-Furthermore, for classification tasks on MNIST and Fashion-MNIST, our approach demonstrates superior classification accuracy compared to the other benchmarks, shown in Tab.1.
-This enhancement is consistent with theoretical expectations, as our methodology learns the inherent probabilistic structure of the training data rather than depending on predetermined heuristics, static partitioning, or hash-based approximations.
-By dynamically adapting to the statistical properties of the data distribution, our method identifies more semantically coherent neighbors.
-Our results suggest that GARLIC acts as a local manifold modeler, not just an indexing method offering smoother, distribution-aware retrieval and classification.
-Table 2: Effect of loss terms and adaptive refinement on average Recall10@10 / Probe (×$\times 10^{5}$). Higher is better.
-Configuration Performance $\uparrow$
-Loss terms
-   w/ all 10.43
-   $\mathcal{L}_{div}+\mathcal{L}_{cov}$ 9.82
-   $\mathcal{L}_{div}+\mathcal{L}_{anchor}$ 7.87
-   $\mathcal{L}_{div}$ 7.17
-[…]
-   w/ both 10.43
-   w/ split 8.65
-   w/ clone 8.47
-   w/o any 2.79
-Ablation study.
-We investigate the robustness and data efficiency of our method through controlled downsampling of the training set.
-As shown in Fig.
-4(a), our method consistently maintains strong performance across varying training sizes.
-Notably, even with only $1\%$ of the training data, our model outperforms the Faiss-IVF variants of up to $10\%$ training size in terms of Recall@1 versus the number of retrieved candidates.
-This trend remains consistent across higher percentages, demonstrating that our approach learns a compact yet highly effective representation of the data distribution.
-Furthermore, Fig.
-4(b) highlights the favorable scaling of our method with respect to training time, while Fig.
-4(c) demonstrates the impact of our quantization scheme.
-In particular, it exhibits strong performance across a broad quantization range ($20-100\%$), with only the most aggressive setting (10%) leading to degradation.
-This indicates that our quantization strategy is robust and well-aligned with our model structure.
-Fig.
-4(c) shows that increasing the number of initial Gaussians accelerates convergence.
-However, the primary role of adaptive refinement is not to fix the final count, but to control their spatial density.
-Tab.
-2 analyzes the impact of loss terms and adaptive refinement.
-Split and clone are pivotal for performance, while the covariance term ($\mathcal{L}_{\text{cov}}$) significantly boosts retrieval.
-The anchor term ($\mathcal{L}_{\text{anchor}}$) has minor effect on raw performance but is essential for making Gaussians geometrically informative.
-Additional ablations and results, covering hyperparameter sensitivity, design variants, and the ability of Gaussians to capture non-linear, informative structure, are provided in the SM.
-Limitations.
-GARLIC uses Mahalanobis distance, which assumes Euclidean structure and is not directly compatible with angular similarity.
-This can be addressed by replacing Gaussians with angular analogues such as von Mises–Fisher distributions or by embedding into hyperspherical spaces.
-Although GARLIC performs strongly in low-probe regimes, achieving very high recall (e.g., $\geq 95\%$) requires more computation.
-This recall-latency trade-off can be mitigated by increasing the number of Gaussians and structuring them with trees (e.g., KD-tree or Ball-tree over Gaussian means), reducing query cost.
-The space complexity scales quadratically with the dimension ($O(Kd^{2})$) due to full anisotropic covariances; however, this can be alleviated by using diagonal or low-rank approximations, or by quantizing the Cholesky factors.
-Lastly, while the method is mildly sensitive to initialization, its impact is reduced by the information-theoretic objectives and the progressive refinement through split and clone operations.
-4 Conclusions
-We introduced GARLIC, a geometric structure that learns the underlying distribution for both approximate nearest neighbor search and classification.
-By combining information-theoretic objectives with adaptive refinement (split and clone), and representing the space via anisotropic Gaussians with local quantization, our method achieves strong performance in both retrieval and classification, particularly in low-probe regimes.
-Experiments demonstrate competitive recall-efficiency tradeoffs and robustness under severe data reduction, highlighting its generalization capabilities.
-Acknowledgments
-PR and IE were partially supported by project MIS 5154714 of the National Recovery and Resilience Plan Greece 2.0 funded by the European Union under the NextGenerationEU Program.
-PR, CT, IC, GI and IE have been partially supported by the European Union and the “Greece 2.0” national recovery and resilience plan, under the call "RESEARCH-CREATE-INNOVATE" (Code: ΤΑΕΔΚ-06168).
-Appendix A Supplementary Material
-This supplementary material provides additional details and experiments as mentioned in the main paper.
-Sec.
-A.1 discusses and provides detailed computational time and space complexity for the build and query procedures of our proposed method.
-Sec.
-A.2 contains additional technical details regarding the datasets used and the experimental setup, as well as the training pipeline.
-Sec.
-A.3 provides ablation studies to validate our choices regarding the tuned hyperparameters of our method and the components of it (discussed in Sec.4 of the main paper).
-Sec.
-A.4 contains visualization of diagnostics and statistics regarding our proposed method, in order to provide a comprehensive understanding of its behavior.
-A.1 Complexity Analysis
-We analyze the computational time and space complexity of our method in three parts: index construction, query execution, and storage.
-The analysis is expressed in terms of standard parameters, including the dataset size $|X|$, embedding dimension $d$, the number of Gaussians $K$, and the reduced PCA dimension $r\ll d$.
-Our goal is to ensure that each component remains scalable with respect to high-dimensional data and large-scale datasets.
-We summarize the complexity of each phase below.
-Index build complexity.
-Let $|X|$ be the number of data points, $d$ the embedding dimension, $K^{\prime}$ the initial number of Gaussians, and $r\ll d$ the reduced dimension used for local quantization.
-Let $I$ be the number of optimization steps.
-We denote by $S$, $C$, and $P$ the number of splits, clones, and pruned Gaussians, respectively, and define the final number of Gaussians as $K=K^{\prime}+S+C-P$.
-Let $|B_{g}|$ be the average bucket size, $c$ the number of K-Means clusters used during splitting, and $k^{\prime}$ the number of candidate points sampled per clone.
-For the initialization, since we use K-Means++ on $K^{\prime}$ total cluster centers, we need $\mathcal{O}(K^{\prime}\cdot d\cdot|X|)$ time.
-For the optimization part, we need to perform a total number of $I$ iterations of full Mahalanobis-based point-to-Gaussian assignment, thus a total of $\mathcal{O}(I\cdot|X|\cdot K\cdot d^{2})$ worst-case time.
-Separate from the optimization, we analyze the split, clone and prune operations that are not applied on every iteration of the optimization.
-(i) The split operation runs DBSCAN or K-Means (with $c$ clusters), thus for a total of $S$ such operations we would need $\mathcal{O}(S\cdot(|B_{g}|\cdot d\cdot c+d^{2}))$; (ii) the clone operation locates the subset of points outside the Gaussian’s boundary (e.g., between $\tau\cdot\sigma$ and $(1+\epsilon)\cdot\tau\cdot\sigma$) for which it identifies new local modes, thus for a total of $C$ operations, it leads to $\mathcal{O}(C\cdot(k^{\prime 2}\cdot d+d^{2}))$; and (iii) the prune operation simply removes low-cardinality Gaussians and reassigns the points to the nearest active Gaussian, which takes $\mathcal{O}(P\cdot|B_{g}|\cdot d)$ time.
-For the quantization of each Gaussian, PCA is performed on all points inside the Gaussian, which projects the data into reduced local bases.
-In total, for the quantization we need $\mathcal{O}(K^{\prime}\cdot|B_{g}|\cdot d^{2}+|X|\cdot d\cdot r)$ time.
-Query complexity.
-Let $K$ be the number of Gaussians, $k$ the number selected per query, $d$ the dimension, $r$ the PCA dimension, $b$ the number of bins per Gaussian, $T$ the number of optimization steps to find the shortest distance from the query point to the boundary of a spherical bin in the reduced PCA space, $\rho$ the probed bin ratio, and $\beta$ the average bin size.
-For a single query, we first need to measure distances from the set of Gaussians, which takes $\mathcal{O}(K\cdot d^{2})$.
-Then, for the $k$ nearest Gaussians we need to locate the subset of data to be examined.
-For this, for each of the selected $k$ Gaussians, we need to compute the local PCA projections ($\mathcal{O}(k\cdot d\cdot r)$), then compute and sort the spherical distances to all $b$ bins ($\mathcal{O}(k\cdot b\cdot r\cdot T)$), of which only the $\rho$ fraction is probed.
-From each, up to $\beta$ candidates are gathered and re-ranked using Euclidean distance, which needs $\mathcal{O}(k\cdot\rho\cdot b\cdot\beta\cdot d)$ time.
-In practice, the re-ranking factor dominates the complexity, which is sublinear.
-Space complexity.
-Let $K$ be the number of Gaussians, $d$ the data dimension, and $N=|X|$ the dataset size.
-The model stores mean vectors $\boldsymbol{\mu}\in\mathbb{R}^{K\times d}$, Cholesky parameters $\mathbf{L}\in\mathbb{R}^{K\times d\times d}$, and Buckets storing point indices, requiring $\mathcal{O}(N)$ space.
-Thus, the total space complexity is:
-$\mathcal{O}\left(K\cdot(d^{2}+d)+N\right)$
-where $K\cdot d^{2}$ dominates.
-Still, space complexity can be reduced to $\mathcal{O}\left(K\cdot d\right)$ by enforcing diagonal covariance matrices, at the expense of reduced expressiveness in anisotropic regions of the space.
-A.2 Implementation details
-This subsection summarizes all implementation and training-specific parameters used in our model, including optimizer schedules, architectural constants, and adaptive procedures such as splitting, cloning and pruning.
-These details provide context for reproducibility and support the complexity analysis in the main paper.
-Furthermore, there are dataset specifics, such as licenses and descriptions.
-Table 3: Dataset Information
-Dataset Dimension Size Description License
-SIFT1M 128 1M Image descriptors CC0
-MNIST 784 70K Handwritten digits CC BY-SA 3.0
-Fashion-MNIST 784 70K Fashion items MIT
-Datasets.
-We evaluate our method on three standard benchmark datasets: (1) SIFT1M [37], containing one million 128-dimensional SIFT descriptors that capture scale- and rotation- invariant local image features; (2) MNIST [32], consisting of 70,000 grayscale handwritten digit images (28×28 pixels) flattened to 784-dimensional vectors; and (3) Fashion-MNIST [44], a more challenging variant with the same format but featuring 10 categories of fashion items.
-For retrieval tasks, we used the ANN-Benchmark [8] versions of these datasets (available at http://ann-benchmarks.com/sift-128-euclidean.hdf5, http://ann-benchmarks.com/mnist-784-euclidean.hdf5, and http://ann-benchmarks.com/fashion-mnist-784-euclidean.hdf5) to ensure a standardized comparison with existing methods.
-For MNIST and Fashion-MNIST classification experiments, we utilized the PyTorch torchvision implementations to maintain consistency with established classification benchmarks.
-Experimental setup.
-Our method was implemented in Python, using optimized libraries such as PyTorch [40] and NumPy [22].
-The experiments were carried out on an Intel Core i7-7820X CPU (16 threads), a Quadro RTX 8000 GPU (48 GB VRAM) and 125 GiB of RAM.
-Training configuration.
-Batch sizes of $20000$ are used for SIFT, and $5000$ for MNIST and Fashion-MNIST datasets, with normalization by subtracting the mean and dividing by the standard deviation.
-The number of training epochs is $250$, with a typical early stop at $120$.
-Gaussian updates are scheduled with two phases: a warm-up phase lasting $35$ epochs and an optimization phase where structural operations like splitting and cloning are triggered every $35$ epochs and pruning every $60$.
-Learning rate schedule.
-The learning rates follow a linear warm-up and exponential decay scheme.
-Specifically, for the Cholesky parameters, the rates are subjected to a warm-up phase from $1\times 10^{-7}$ to $5\times 10^{-4}$, followed by a decay to $9\times 10^{-5}$.
-In terms of the means, they are warmed up from $1\times 10^{-7}$ to $9\times 10^{-3}$, and subsequently decay to $3\times 10^{-3}$.
-Notably, the learning rate associated with the means is maintained at a relatively higher level than that of the covariances.
-This approach is designed to promote the adjustment of the Gaussian centers rather than the expansion of their radii.
-Adaptive refinement.
-Splitting is applied to Gaussians with cardinality exceeding a fraction of the dataset $\gamma=1\times 10^{-2}$, using DBSCAN or K-Means, as fallback, with $c=2$ clusters.
-The covariance of each new Gaussian is scaled down by $9\times 10^{-1}$.
-Cloning selects dense regions just outside the Gaussian boundary, defined by a Mahalanobis shell with inner threshold $\tau$ and outer threshold $(1+\epsilon)\cdot\tau$, where $\epsilon=2.2$ controls the shell thickness.
-From the set of points referenced in $(\tau,e\tau]$, a random selection of $60\%$ is made.
-Cloning is not performed on a Gaussian unless its cardinality surpasses a threshold specified by $8\times 10^{-4}|X|$.
-Gaussians that have degenerated into a single point are eliminated.
-Pruning is executed at intervals of every $60$ epochs.
-Quantization.
-Local PCA is performed per bucket using top-$3$ eigenvectors.
-Reduced points are quantized using a spherical grid with $n_{\text{radial}}=6$ and $n_{\text{angular}}=4$, forming directional bins per Gaussian.
-Loss.
-The total loss is a weighted sum of three components: divergence $\lambda_{div}=1.0$, covariance $\lambda_{cov}=1.0$, and anchor term $\lambda_{anchor}=10^{-2}$, with a weight $\alpha=10^{-1}$ that balances position and shape.
-When calculating the $\mathcal{L}_{cov}$ loss, a numerical epsilon of $1\times 10^{-12}$ is used to ensure stability.
-A.3 Additional Ablation Studies
-[…]
+Across datasets we observe consistent trends: $\varepsilon$-Recall at $1\%$ and $10\%$ rapidly approaches $1.0$, indicating that even under small probe budgets the vast majority of retrieved neighbors lie within $1\%$$10\%$ of the true nearest-neighbor distance.
+The $r95$ and $r99$ curves remain close to $1.0$, showing that $5\%99$$9\%$ of queries admit near-exact retrieval, while $RE_{\text{mean}}$ stays low and stable, confirming that average distortion is minimal.
+Together with high P@10, these results demonstrate that the learned Gaussian partitions not only achieve strong recall but also return candidates that are quantitatively close to the exact nearest neighbors, providing both efficiency and fidelity in the ANN process.
+Additional Abblation Studies
+In Figure 7(7(a)), the contribution of each loss term to the optimization process is analyzed.
+The results indicate that the covariance loss $\mathcal{L}_{\text{cov}}$ produces the most significant performance improvement, while the anchor loss $\mathcal{L}_{\text{anchor}}$ functions as an effective regularizer, grounding each Gaussian by aligning it with its corresponding point distribution.
+Figure 7(7(b)) demonstrates the effects of our adaptive refinement operations, where the split and clone mechanisms improve retrieval performance when used together, each contributing by providing complementary benefits to the quality of representation.
+Figure 7(7(c)), addresses the parameter $\tau$ within $\mathcal{L}_{div}$, demonstrating that $\tau=3$ is the optimal choice, particularly in scenarios involving a limited number of probes.
+Values surpassing those presented in our study ($\tau>3$) result in less optimal outcomes.
+For instance, $\tau=4$ demonstrates a Recall@1 value of 0.67~0.67 when evaluated with $154264$ probes.
 (a) Loss terms
@@
 (c) ReLU parameters $\tau$
-(d) High-recall
+(d) Quantization
+(e) Dimensionality effect
 […]
-Architectural components.
-In Fig.
-5(5(a)), the contribution of each loss term to the optimization process is analyzed.
-The results indicate that the covariance loss $\mathcal{L}_{\text{cov}}$ produces the most significant performance improvement, while the anchor loss $\mathcal{L}_{\text{anchor}}$ functions as an effective regularizer, grounding each Gaussian by aligning it with its corresponding point distribution.
-Fig.
-5(5(b)) demonstrates the effects of our adaptive refinement operations, where the split and clone mechanisms improve retrieval performance when used together, each contributing by providing complementary benefits to the quality of representation.
-Fig.
-5(5(c)), addresses the parameter $\tau$ within $\mathcal{L}_{div}$, demonstrating that $\tau=3$ is the optimal choice, particularly in scenarios involving a limited number of probes.
-Values surpassing those presented in our study ($\tau>3$) result in less optimal outcomes.
-For instance, $\tau=4$ demonstrates a Recall10@10 value of 0.670.67~{}0.670.67 when evaluated with $65000$ probes.
-A way to address the limitation associated with high-recall scenarios, as discussed in the limitations paragraph of the main paper, is show in Fig.
-5(5(d)).
-Specifically, it involves assessing all Gaussian candidates that meet the condition $d_{M}(x,g)\leq\tau$, rather than solely the one that minimizes the Mahalanobis distance.
-This approach is also semantically justified, considering that multiple Gaussians may occupy the same region.
-Consequently, a hybrid approach can be used, utilizing the $\arg\min$ strategy for scenarios with a low number of probes and employing all candidates that satisfy condition $d_{M}(x,g)\leq\tau$ for high-recall cases, thus effectively addressing the limitation.
-(a) Initialization
-(b) Split threshold $\gamma$
-(c) Clone threshold $e$
-(d) Dimensionality effect
-[…]
-Table 4: Effect of anisotropy on average Recall10@10 / Probe (×$\times 10^{5}$). Higher is better. As anisotropy decreases, performance degrades due to excessive probe usage.
+Furthermore, Figure 7(7(d)) demonstrates the impact of our quantization scheme.
+In particular, it exhibits strong performance across a broad quantization range ($20{-}100\%$), with only the most aggressive setting (10%) leading to degradation.
+This indicates that our quantization strategy is robust and well-aligned with our model structure.
+Finally, Figure 7(7(e)) examines dimensionality effects using Fashion-MNIST data resized to various dimensions (to simulate higher-dimensional embeddings).
+The results confirm that GARLIC maintains strong performance across a wide range of dimensionalities.
+Table 2: Effect of anisotropy on average Recall@1 / Probe ($\times 10^{5}$). Higher is better. As anisotropy decreases, performance degrades due to excessive probe usage.
 Configuration Performance $\uparrow$
 Covariance structure
-   Anisotropic 10.43
-   Diagonal 11.04
-   Isotropic 10.63
-Validation of hyperparameters.
-Fig.
-6 provides a systematic evaluation of our key hyperparameters.
-In Fig.
-6(6(a)), we investigate initialization strategies for Gaussian centers, revealing that K-Means initialization consistently outperforms alternatives.
-Although farthest point sampling (FPS) achieves comparable recall, it requires approximately 50% more probes, and catching up to K-Means requires more training time.
-In Fig.
-6(6(b)), the split threshold $\gamma$, controlling the density condition for Gaussian subdivision, is examined.
-The optimal values are $\gamma=0.005$ and $\gamma=0.01$, though $\gamma=0.005$ requires more time due to more frequent splits.
-As for the clone threshold $e$, depicted in Fig.
-6(6(c)), which determines the outer shell for cloning, it remains stable across different values, with $e=2.2$ being preferable in situations requiring high recall.
-Finally, Fig.
-6(6(d)) examines dimensionality effects using Fashion-MNIST data resized to various dimensions.
-The results confirm that GARLIC maintains strong performance across a wide range of dimensionalities, highlighting its robustness to the feature space complexity.
+   Anisotropic 16.20
+   Diagonal 11.22
+   Isotropic 10.75
 Covariance structure vs performance.
-We conduct a targeted study to assess the impact of different covariance configurations-namely, full (anisotropic), diagonal, and isotropic-baffect the performance of GARLIC.
-As shown in Table 4, reducing the Gaussian expressiveness from full to diagonal and then to isotropic leads to a notable decline in average Recall10@10 per probe.
+We conduct a targeted study to assess the impact of different covariance configurations-namely, full (anisotropic), diagonal, and isotropic-affect the performance of GARLIC.
+As shown in Table 2, reducing the Gaussian expressiveness from full to diagonal and then to isotropic leads to a notable decline in average Recall@1 per probe.
 While diagonal and isotropic configurations reduce both the parameter count and computation per Gaussian to a linear level, $\mathcal{O}(k\cdot d)$, they lead to aggressive pruning and an increased number of probes to make up for the representational loss.
 This suggests that space complexity cannot be drastically reduced without harming retrieval quality because simpler Gaussian parameterizations result in degraded locality and coverage.
-A.4 Qualitative Results
+A.1.1 Qualitative Results
 To gain a more comprehensive understanding of the model’s behavior, we present a collection of diagnostic visualizations applied to the Fashion-MNIST, MNIST, and SIFT datasets.
@@
 Curvature-based diagnostics.
-Fig.
-8 provides four views exploring the relationship between curvature and structural properties of the Gaussian assignments across datasets, to examine wheather the learned model is informative and geometrically consistent with the data $x\in X$.
+Figure 9 provides four views exploring the relationship between curvature and structural properties of the Gaussian assignments across datasets, to examine wheather the learned model is informative and geometrically consistent with the data $x\in X$.
 The first row shows the standard deviation of local curvature as a function of the average radius ($l_{2}$) per Gaussian.
@@
 Each subplot shows a different measure: anisotropy, band energy, diagonality, effective dimension, and explained variance.
-Fig.
-9 demonstrates the capability of GARLIC to capture the intrinsic geometric structure of high-dimensional data with locally varying dimensionality.
+Figure 10 demonstrates the capability of GARLIC to capture the intrinsic geometric structure of high-dimensional data with locally varying dimensionality.
 The anisotropy histograms (first row) reveal how Gaussians adapt to regions of different local dimensionality, with $\lambda_{\text{max}}/\lambda_{\text{min}}$ ratios ranging from nearly isotropic to highly stretched configurations across all datasets.
@@
 This completes the supplementary material.
+A.2 Implementation details
+This subsection summarizes all implementation and training-specific parameters used in our model, including optimizer schedules, architectural constants, and adaptive procedures such as splitting, cloning and pruning.
+These details provide context for reproducibility and support the complexity analysis in the main paper.
+Furthermore, there are dataset specifics, such as licenses and descriptions.
+Table 3: Dataset Information
+Dataset Dimension Size Description License
+SIFT1M 128 1M Image descriptors CC0
+MNIST 784 70K Handwritten digits CC BY-SA 3.0
+Fashion-MNIST 784 70K Fashion items MIT
+Datasets.
+We evaluate our method on three standard benchmark datasets: (1) SIFT1M Lowe (2004), containing one million 128-dimensional SIFT descriptors that capture scale- and rotation- invariant local image features; (2) MNIST LeCun et al.
+(1998), consisting of 70,000 grayscale handwritten digit images (28×28 pixels) flattened to 784-dimensional vectors; and (3) Fashion-MNIST Xiao et al.
+(2017), a more challenging variant with the same format but featuring 10 categories of fashion items.
+For retrieval tasks, we used the ANN-Benchmark Aumüller et al.
+(2020) versions of these datasets (available at http://ann-benchmarks.com/sift-128-euclidean.hdf5, http://ann-benchmarks.com/mnist-784-euclidean.hdf5, and http://ann-benchmarks.com/fashion-mnist-784-euclidean.hdf5) to ensure a standardized comparison with existing methods.
+Experimental setup.
+Our method was implemented in Python, using optimized libraries such as PyTorch Paszke (2019) and NumPy Harris et al.
+(2020).
+The experiments were carried out on an Intel Core i7-7820X CPU (16 threads), a Quadro RTX 8000 GPU (48 GB VRAM) and 125 GiB of RAM.
+Training configuration.
+Batch sizes of $50000$ are used for SIFT, and $20000$ for MNIST and Fashion-MNIST datasets, with z-score normalization by subtracting the mean and dividing by the standard deviation.
+The number of training epochs is $250$, with a typical early stop at $120$.
+Gaussian updates are scheduled with two phases: a warm-up phase lasting $35$ epochs and an optimization phase where structural operations like splitting and cloning are triggered every $35$ epochs and pruning every $60$.
+Learning rate schedule.
+The learning rates follow a linear warm-up and exponential decay scheme.
+Specifically, for the Cholesky parameters, the rates are subjected to a warm-up phase from $1\times 10^{-7}$ to $5\times 10^{-4}$, followed by a decay to $9\times 10^{-5}$.
+In terms of the means, they are warmed up from $1\times 10^{-7}$ to $9\times 10^{-3}$, and subsequently decay to $3\times 10^{-3}$.
+Notably, the learning rate associated with the means is maintained at a relatively higher level than that of the covariances.
+This approach is designed to promote the adjustment of the Gaussian centers rather than the expansion of their radii.
+Adaptive refinement.
+Splitting is applied to Gaussians with cardinality exceeding a fraction of the dataset $\gamma=1\times 10^{-2}$, using DBSCAN or K-Means, as fallback, with $c=2$ clusters.
+The covariance of each new Gaussian is scaled down by $9\times 10^{-1}$.
+Cloning selects dense regions just outside the Gaussian boundary, defined by a Mahalanobis shell with inner threshold $\tau$ and outer threshold $(1+e)\cdot\tau$, where $e=2.2$ controls the shell thickness.
+From the set of points referenced in $(\tau,e\tau]$, a random selection of $60\%$ is made.
+Cloning is not performed on a Gaussian unless its cardinality surpasses a threshold specified by $8\times 10^{-4}|X|$.
+Gaussians that have degenerated into a single point are eliminated.
+Pruning is executed at intervals of every $60$ epochs.
+Quantization.
+Local PCA is performed per cell using top-$3$ eigenvectors.
+Reduced points are quantized using a spherical grid with $n_{\text{radial}}=6$ and $n_{\text{angular}}=4$, forming directional bins per Gaussian.
+Loss.
+The total loss is a weighted sum of three components: divergence $\lambda_{div}=1.0$, covariance $\lambda_{cov}=1.0$, and anchor term $\lambda_{anchor}=10^{-2}$, with a weight $\alpha=10^{-1}$ that balances position and shape.
+When calculating the $\mathcal{L}_{cov}$ loss, a numerical epsilon of $1\times 10^{-12}$ is used to ensure stability.
+A.3 Complexity Analysis
+We analyze the computational time and space complexity of our method in three parts: index construction, query execution, and storage.
+The analysis is expressed in terms of standard parameters, including the dataset size $|X|$, embedding dimension $d$, the number of Gaussians $K$, and the reduced PCA dimension $r\ll d$.
+Our goal is to ensure that each component remains scalable with respect to high-dimensional data and large-scale datasets.
+We summarize the complexity of each phase below.
+Index build complexity.
+Let $I$ be the number of optimization steps and $K^{\prime}$ the initial number of Gaussians.
+We denote by $S$, $C$, and $P$ the number of splits, clones, and pruned Gaussians, respectively, and define the final number of Gaussians as $K=K^{\prime}+S+C-P$.
+Let $|B_{g}|$ be the average cell size, $c$ the number of K-Means clusters used during splitting, and $k^{\prime}$ the number of candidate points sampled per cloning operation.
+For the initialization, since we use K-Means++ on $K^{\prime}$ total cluster centers, we need $\mathcal{O}(K^{\prime}\cdot d\cdot|X|)$ time.
+For the optimization part, we need to perform a total number of $I$ iterations of full Mahalanobis-based point-to-Gaussian assignment, thus a total of $\mathcal{O}(I\cdot|X|\cdot K\cdot d^{2})$ worst-case time.
+Separate from the optimization, we analyze the split, clone and prune operations that are not applied on every iteration of the optimization.
+(i) The split operation runs DBSCAN or K-Means (with $c$ clusters), thus for a total of $S$ such operations we would need $\mathcal{O}(S\cdot(|B_{g}|\cdot d\cdot c+d^{2}))$; (ii) the clone operation locates the subset of points outside the Gaussian’s boundary (between $\tau\cdot\sigma$ and $(1+\epsilon)\cdot\tau\cdot\sigma$) for which it identifies new local modes, thus for a total of $C$ operations, it leads to $\mathcal{O}(C\cdot(k^{\prime 2}\cdot d+d^{2}))$; and (iii) the prune operation simply removes low-cardinality Gaussians and reassigns the points to the nearest active Gaussian, which takes $\mathcal{O}(P\cdot|B_{g}|\cdot d)$ time.
+For the quantization of each Gaussian, PCA is performed on all points inside the Gaussian, which projects the data into reduced local bases.
+In total, for the quantization we need $\mathcal{O}(K^{\prime}\cdot|B_{g}|\cdot d^{2}+|X|\cdot d\cdot r)$ time.
+From all the terms described, the optimization term dominates.
+Query complexity.
+Let $K$ be the number of Gaussians, $k$ the number selected per query, $d$ the dimension, $r$ the PCA dimension, $b$ the number of bins per Gaussian, $T$ the number of optimization steps to find the shortest distance from the query point to the boundary of a spherical bin in the reduced PCA space, $\rho$ the probed bin ratio, and $\beta$ the average bin size.
+For a single query, we first need to measure distances from the set of Gaussians, which takes $\mathcal{O}(K\cdot d^{2})$.
+Then, for the $k$ nearest Gaussians we need to locate the subset of data to be examined.
+For this, for each of the selected $k$ Gaussians, we need to compute the local PCA projections ($\mathcal{O}(k\cdot d\cdot r)$), then compute and sort the spherical distances to all $b$ bins ($\mathcal{O}(k\cdot b\cdot r\cdot T)$), of which only the $\rho$ fraction is probed.
+From each, up to $\beta$ candidates are gathered and re-ranked using Euclidean distance, which needs $\mathcal{O}(k\cdot\rho\cdot b\cdot\beta\cdot d)$ time.
+In practice, the re-ranking factor dominates the complexity, which is sublinear.
+Space complexity.
+Let $K$ be the number of Gaussians, $d$ the data dimension, and $N=|X|$ the dataset size.
+The model stores mean vectors $\boldsymbol{\mu}\in\mathbb{R}^{K\times d}$, Cholesky parameters $\mathbf{L}\in\mathbb{R}^{K\times d\times d}$, and cells storing point indices, requiring $\mathcal{O}(N)$ space.
+Thus, the total space complexity is:
+$\mathcal{O}\left(K\cdot(d^{2}+d)+N\right)$
+where $K\cdot d^{2}$ dominates.
+Still, space complexity can be reduced to $\mathcal{O}\left(K\cdot d\right)$ by enforcing diagonal covariance matrices, at the expense of reduced expressiveness in anisotropic regions of the space.